### v1

# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

Abstract
Text-guided diffusion models catalyze a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts. Recent advancements introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pre-trained diffusion models for audio modification. Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy. And the lack of attention control hinders the fine-grained manipulations of music. To counteract these limitations, we introduce the Disentangled Inversion technique, which is designed to disentangle the diffusion process into triple branches, thereby magnifying their individual capabilities for both precise editing and preservation. Furthermore, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention and cross-attention with an additional Harmonic Branch to achieve the desired composition and structural information in the target music. Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing whilst safeguarding structural integrity. To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across 10 distinct editing categories, which facilitates both zero-shot and instruction-based music editing tasks. Our method demonstrates unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques. [^1] Audio samples are available at https://MEDIC-Zero.github.io/. [^2] Benchmark data are available at https://huggingface.co/datasets/liuhuadai/ZoME-Bench.
1 Introduction
Advancements in text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have markedly enhanced audio generation, leveraging their impressive capability for realistic and varied outputs. These models (Liu et al., 2023b; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes in response to specific textual prompts. Early audio editing strategies rely on training models from scratch or test-time optimization, both of which are hampered by intensive computational demands. Recent advances (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through DDIM (Song et al., 2020) and DDPM (Ho et al., ) inversion techniques, but challenges remain.
Key among these challenges is maintaining editing fidelity - ensuring that the model’s output matches the given instructions - and preserving essential content, a hallmark of effective audio editing that targets only necessary changes. Balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving suboptimal for conditional diffusions (Mokady et al., 2023). Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space, yet at the expense of reduced modification capabilities due to noise reduction.
In this work, we rigorously examine the shortcomings of the DDIM inversion approach. Our comprehensive analysis, suggests that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence. The implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) further amplifies the accumulated errors. Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has demonstrated its power in achieving high fidelity and essential content preservation. Yet, analogous control mechanisms in music editing are noticeably absent. To bridge this gap, we introduce an innovative Disentangled Inversion Control technique. This technique is comprised of two main components: Harmonic Attention Control and Disentangled Inversion. The former integrates cross-attention and mutual self-attention controls, supplemented by a novel Harmonic Branch, which facilitates precise, fine-grained manipulations for both rigid and non-rigid editing scenarios. The latter component strategically deconstructs the diffusion process into three distinct branches, enabling each to realize its full potential independently.
Acknowledging the lack of standardized benchmarks in the field of audio editing, we introduce ZoME-Bench. The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories. Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing. Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
Our contributions can be summarized as follows. 1) We introduce a novel, tuning-free methodology referred to as Disentangled Inversion Control (DIC), designed to facilitate both consistent manipulations of musical elements and intricate non-rigid editing tasks. 2) An Harmonic Attention Control framework is introduced to unify cross-attention control and self-attention control, which enables both rigid and non-rigid editing. 3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error without optimization by branch disentanglement, aiding in accurately editing the music while preserving the structural information. 4) We introduce a tailored benchmark for music editing, named ZoME-Bench, and Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
2 Related Works
2.1 Text-based Audio editing
The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt. Two primary challenges command intense focus. The first is the fidelity of editing, which pertains to the model’s ability to strictly follow the initial instructions. The second challenge involves essential content preservation, a critical attribute of competent audio editing models that ensures only suitable regional modifications are made during the editing process. Existing methodologies for addressing these intricate challenges typically follow one of three paths. The first involves attempts to develop end-to-end editing models (Copet et al., 2023; Agostinelli et al., 2023; Chen et al., 2024)that employ diffusion processes. However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets. The second path involves test-time optimization strategies that utilize large pre-trained models for editing (Paissan et al., 2023; Plitsis et al., 2024). Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text-embeddings for signal reconstruction. Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load. The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020) inversion strategies to extract diffusion noise vectors that match the source signal. Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework. In this study, we propose a new inversion technique named Disentangled Inversion Control. This technique aims to achieve accurate editing while preserving structural information.
2.2 Inversion Techniques
The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021). While the DDIM inversion proves effective for unconditional diffusion models (Song et al., 2020; Wallace et al., 2023) , its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing. A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges. For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during the editing process. Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024).
Recently, optimization-based inversion methods using specific latent variables have gained popularity (Ju et al., 2023; Kawar et al., 2023). These are designed to minimize accumulated errors stemming from the DDIM inversion. Techniques such as Null-Text Inversion (Mokady et al., 2023) are promising, but they introduce complexity and instability into the optimization process, making it relatively time-consuming. Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
Figure 1: The framework and rationale behind Disentangled Inversion Control. Traditional DDIM inversion introduces incremental errors at each step, leading to only a coarse approximation of the source audio’s distribution. The implementation of classifier-free guidance exacerbates the error accumulation. In contrast, our method intricately separates the diffusion process into a triple-branch system, allowing each branch to fully realize its individual potential.
3 Disentangled Inversion Control
3.1 Motivation
Figure 1 and Preliminaries in Appendix 2 reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence. The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
In the landscape of prompt-based editing (Dong et al., 2023; Kim et al., 2022; Feng et al., 2023), the ability to grasp the subtleties of linguistics and enable more granular cross-modal interactions stands as a formidable challenge.  Hertz et al. (2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$. This sparks a plethora of attention control techniques that direct the image domain’s target denoiser network $\hat{\epsilon}_{\theta}$ to better reflect target prompts. Yet, analogous control mechanisms in music editing are noticeably absent.
Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach poised to achieve fine-grained zero-shot music editing. DIC strategically disentangles the diffusion process as a triple-branch, granting each branch to optimize its functionality. At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance. We will first introduce Harmonized Attention Control in Section 3.2 and discuss Disentangled Inversion in Section 3.3.
3.2 Harmonized Attention Control Framework
The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023a). At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer. The attention mechanism is formalized as:
$\text{Attention}(Q,K,V)=MV=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (1)
where $Q$ denotes the query features derived from the audio features, $K$, $V$ represent the key and value features, and $M$ is the attention maps. We explore varying semantic transformations of audio content through bifurcated attention control strategies—cross-attention for rigid changes and mutual self-attention for non-rigid adjustments. Finally, we introduce an intermediate branch to host the desired composition and structural information in the target music.
Figure 2: The framework of Harmonized attention control (HAC). HAC unifies cross-attention control and mutual self-attention control with an additional branch named Harmonic Branch to host the desired composition and structural information in the target music.
3.2.1 Cross-Attention Control
Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the modified prompt ${\mathcal{P}}^{*}$. We follow the practice of Prompt-to-Prompt and define CAC as Global Attention Refinement and Local Attention Blend.
Global Attention Refinement
At a given time step $t$, the attention map $M_{t}$ for both the origin and target branches is computed, averaging over all layers with respect to the noised latent $\mathbf{z}_{t}$. We employ an alignment function $A$ that maps each token index from the target prompt $\mathcal{P}^{*}$ to its equivalent in $\mathcal{P}$ or to None for non-aligning tokens. The refinement action is thus:
$\text{Refine}(M^{src}_{t},M^{tgt}_{t},t)=\begin{cases}(M^{tgt}_{t})_{i,j}&\text{if }A(j)=\text{None},\\ (M_{t})_{i,j}&\text{otherwise}.\end{cases}$ (2)
Local Attention Blends
Beyond global attention, we incorporate a blending mechanism as suggested by  Hertz et al. (2022) and  Mokady et al. (2023). This method selectively integrates and maintains certain semantics by using target blend words $w^{tgt}$ for semantic additions and source blend words $w^{src}$ for semantic preservation. At each denoising step $t$, the mechanism operates on the target latent ${\bm{z}}^{tgt}_{t}$ as follows:
$m_{\text{tgt}}$ $=\text{Threshold}\left[M^{tgt}_{t},w_{\text{tgt}},k_{\text{tgt}}\right],$ (3)
$m_{\text{src}}$ $=\text{Threshold}\left[M^{src}_{t},w_{\text{src}},k_{\text{src}}\right],$ (4)
${\bm{z}}^{tgt}_{t}$ $=(1-m^{tgt}+m^{src})\odot{\bm{z}}^{src}_{t}+(m^{tgt}-m^{src})\odot{\bm{z}}^{tgt}_{t}$ (5)
where $m_{\text{tgt}}$ and $m_{\text{src}}$ are binary masks and threshold function is as delineated below:
$\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$ (6)
For simplificity, we define the process of local editing in Equation 10 as:
${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$ (7)
Scheduling Cross-Attention Control
Applying cross-attention control only at early stages ensures creative flexibility while maintaining structure. Acting on insights from  Hertz et al. (2022), we limit cross-attention to the initial phases up to a cutoff point $\tau_{c}$. This moderation allows us to capture the nuances and intended changes in musical compositions effectively. The approach is defined as:
$\text{CrossEdit}(M_{t},M^{*}_{t},t)=\begin{cases}\text{Refine}(M^{src},M^{tgt})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
3.2.2 Mutual Self-Attention Control
We diverge from the conventional use of cross-attention mechanisms and instead draw inspiration from the MasaCtrl (Cao et al., 2023) technique to refine music structure through self-attention queries. These queries adeptly navigate through non-rigid musical transformations, aligning with the designated musical theme or instrument (target prompt). The process begins by sketching the foundational musical theme using the target’s self-attention components—$Q^{tgt}$, $K^{tgt}$, and $V^{tgt}$. This is followed by enriching this theme with elements resembling the thematic content from the source ($K^{src}$, $V^{src}$), steered by $Q^{tgt}$. However, applying this attentive modulation uniformly over all processing layers and through every denoising step might result in a composition excessively mirroring the source. Consequently, echoing the ethos of MasaCtrl, our proposed solution selectively employs mutual self-attention in the decoder portion of our music editing U-Net, initiated after a set number of denoising iterations.
Scheduling Mutual Self-Attention Control
The application of mutual self-attention is meticulously planned, beginning at a specific denoising step $S$ and extending beyond a designated layer $L$. The strength and influence of this control mechanism are designed as follows:
$\text{SelfEdit}({Q^{src},K^{src},V^{src}},{Q^{tgt},K^{tgt},V^{tgt}},t)=\begin{cases}{Q^{src},K^{src},V^{src}}&\text{if }t\geq S\text{ and }l\geq L,\\ {Q^{tgt},K^{src},V^{src}}&\text{otherwise}\\ \end{cases}$ (9)
In this framework, $S$ signifies the denoising step from which the mutual self-attention control commences, serving as a temporal threshold. Similarly, $L$ distinguishes the layer index below which this nuanced control strategy becomes operational, tailoring the musical output towards the intended artistic direction.
3.2.3 Harmonic Branch Integretion
The naive combination of cross-attention control and mutual self-attention control sequentially would lead to a sub-optimal outcome in the original dual-branch setup, especially failing the global attention refinement. To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music. Our unified framework is detailed in Algorithm 2 and illustrated in Figure 2. During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$. Following this, cross-attention control is applied on $M_{har}$ and $M_{tgt}$ to refine the semantic information for $M_{tgt}$.
Algorithm 1 Harmonized Attention Control in one DDIM Forward Process
1:  Input: A source prompt ${\mathcal{P}}$, a target prompt ${\mathcal{P}}^{*}$, a source audio latent ${\bm{z}}_{0}$, denoising network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$, current time step $\tau$, source and target blend words $w_{\text{src}},w_{\text{tgt}}$, input latents $z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}$.
2:  $\epsilon_{\text{src}},\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},M_{\text{src}}=\epsilon_{\theta}(z^{\text{src}}_{\tau},\tau,c_{\text{src}})$
3:  $\epsilon_{\text{tgt}},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},M_{\text{tgt}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}})$
4:  $\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\}=\text{SelfEdit}(\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},\tau)$
5:  $\epsilon_{\text{har}},M_{\text{har}}=\epsilon_{\theta}(z^{\text{har}}_{\tau},\tau,c_{\text{src}};\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\})$
6:  $\hat{M^{\text{tgt}}}=\text{CrossEdit}(M_{\text{har}},M_{\text{tgt}},\tau)$
7:  $\hat{\epsilon_{\text{tgt}}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}};\hat{M^{\text{tgt}}})$
8:  $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}=\text{Sample}([z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}],[\epsilon_{\text{src}},\epsilon_{c_{\text{tgt}}},\epsilon_{\text{har}}],\tau)$
9:  ${\bm{z}}^{tgt}_{\tau-1}$ = $\text{LocalEdit}({\bm{z}}^{src}_{\tau-1},{\bm{z}}^{tgt}_{\tau-1},M^{src}_{\tau-1},M^{tgt}_{t},w_{src},w_{tgt})$
10:  Output: $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}$
3.3 Disentangled Inversion Technique
Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal. Increasing the guidance scale enhances editability but sacrifices reconstruction accuracy due to latent code deviation during editing.
In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the target, and the harmonized branch. This bifurcation is designed to optimize the performance of each branch separately. For the source branch, we implement a targeted correction mechanism. By reintegrating the distance $z^{*}_{t}-z^{src}_{t}$ into $z^{src}_{t}$, we directly mitigate the deviation of the pathway. This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio. On the other hand, the target branch and harmonic branch is left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio. The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio. We will further discuss this in the Section 5.4.2. The algorithm of Disentangled Inversion Technique has been outlined in Algorithm 2.
Typical diffusion-based editing (Han et al., 2023b; Miyake et al., 2023) involves two parts: an inversion process to get the diffusion space of the audio, and a forward process to perform editing on the diffusion space. Disentangled Inversion can be plug-and-played into the forward process and rectifies the deviation path step by step. Specifically, Disentangled Inversion first computes the difference between ${\bm{z}}^{*}_{t-1}$ and ${\bm{z}}^{src}_{t-1}$, then adds the difference back to ${\bm{z}}^{src}_{t-1}$ in DDIM forward. We only add the difference of the source prompt in latent space and update ${\bm{z}}^{src}_{t-1}$, which is the key to retaining the editability of the target prompt’s latent space.
Algorithm 2 Disentangled Inversion Technique
1:  Input: A source prompt ${\mathcal{P}}$, a target prompt ${\mathcal{P}}^{*}$, a source audio latent ${\bm{z}}_{0}$, and guidance scale $\omega$.
2:  Output: A edited audio latent ${\bm{z}}^{tgt}_{0}$.
3:  Compute the intermediate results ${\bm{z}}^{*}_{T},...,{\bm{z}}^{*}_{1}$ using DDIM inversion over ${\bm{z}}_{0}$.
4:  Initialize ${\bm{z}}^{src}_{T}\leftarrow{\bm{z}}^{*}_{T}$, ${\bm{z}}^{tgt}_{T}\leftarrow{\bm{z}}^{*}_{T}$, ${\bm{z}}^{har}_{T}\leftarrow{\bm{z}}^{*}_{T}$.
5:  for $t=T$ to $1$ do
6:     $[{\bm{d}}^{src}_{t-1},{\bm{d}}^{tgt}_{t-1},{\bm{d}}^{har}_{t-1}]\leftarrow{\bm{z}}^{*}_{t-1}$ DDIM_Forward$({\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega)$
7:     ${\bm{z}}^{src}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
8:     ${\bm{z}}^{har}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{har}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
9:     ${\bm{z}}^{tgt}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{tgt}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
10:  end for
11:  return  ${\bm{z}}^{tgt}_{0}$
4 BenchMark Construction
Despite recent significant advancements in music generation utilizing the latent diffusion model, music editing has received comparably less attention, especially zero-shot music editing. Hence, to systematically validate our proposed method as a plug-and-play strategy for editing models and compare our method with existing zero-shot music editing methods, as well as compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset, named ZoME-Bench (Zero-shot Music Editing Benchmark). ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps, spanning ten distinct editing types such as add instrument, delete instrument, change genre, and so on. Each sample in this collection is accompanied by its corresponding source prompt, target prompt, human instruction, and the source audio. Additionally, annotations pertinent to attention control methodologies, such as blended words, are meticulously included to facilitate detailed evaluations. We attach the details about our benchmark in Appendix B.
5 Experiments
5.1 Implementation Details
We perform the inference of different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023b) models with 200 inference steps. The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA 4090Ti GPU for a fair comparison. We evaluate all methods in ZoMo-Bench across all editing types. Details can be found in Appendix C.
5.2 Evaluation Metrics
In order to evaluate the performance of our novel Disentangled Inversion Control framework, we employ a diverse set of five metrics that encompass three critical dimensions of audio quality and fidelity: Struture Distance,CLAP Score,LPAPS,MSE (Mean Squared Error)and FAD (Fréchet Audio Distance). All the metrics are introduced in detail in Appendix C.1.
5.3 Zero-shot Music Editing Results
[…]
SDEdit 25.87 0.22 12.18 9.82 0.40
DDPM-Friendly 22.36 0.21 14.08 8.49 0.41
DDIM Inversion 17.11 0.18 6.24 6.12 0.55
DIC 11.97 0.15 2.49 4.54 0.60
Table 1: Comparison with baselines on ZoMo-Bench about struture and content preservation and CLAP similiarity. DIC: our proposed Disentangled Inversion Control.
We present a comparative study of our Disentangled Inversion Control(DIC) against several established music editing techniques, namely SDEdit (Liu et al., ), DDIM Inversion Ho et al. , and DDPM-Friendly (Manor & Michaeli, 2024). Utilizing the ZoMo-Bench test set we released, we assess the quality of the generated audio samples. The assessment primarily concentrates on two key aspects: Essential Content Preservation and Editing Fidelity. The results are summarized in Table 1, leading to two main conclusions:
(1) Our Disentangled Inversion Control demonstrates substantial advancements in maintaining essential content and enhancing editing fidelity, in comparison to the baseline methods. (2) Despite DDIM Inversion achieving high CLAP similarity scores, significantly underperforms in preserving essential content when compared to our approach.
(a) Spider chart for editing quality.
(b) Spider chart for editing consistency.
Figure 3: A comprehensive performance evaluation on the ZoME-bench. We present spider charts of editing quality (CLAP Scores) and consistency (LPAPS) across 9 editing tasks for DDPM-Friendly, DDIM Inversion, and Disentangled Inversion Control methods.
To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing. The LPAPS and CLAP scores are depicted in Figure 3. The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks. (1) While methods like DDIM Inversion exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing fine-grained manipulations, as reflected in their inferior performance, especially in the "Delete Instrument" and "Change Melody" non-rigid editing type.
5.4 Ablation Study
The ablation studies presented in this section aim to validate the contributions of the Unified Attention Control and Refining Dual-path Inversion to our framework’s overall performance. Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in the Appendix D.
5.4.1 Ablation on Attention Control Methods
[…]
HAC 11.97 0.15 2.49 4.54 0.60
w/ CAC 10.24 0.19 2.75 4.76 0.56
w/ MSAC 9.92 0.18 2.50 4.66 0.58
w/ HB 11.75 0.16 2.59 4.67 0.59
Table 2: Ablation results about different control methods. HAC: Harmonized Attention Control.
To ascertain the impact of our attention control mechanisms, we perform ablations on the following configurations: Cross-attention Control Only (w/ CAC), Mutual Self-attention Control only (w/ MSAC), without Harmonic Branch (w/o HB), and any attention control (w/o Control). The findings, detailed in Table 2, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance. (2) The combined use of CAC and MSAC further improves results, but the absence of the harmonic branch yields suboptimal outcomes.
5.4.2 Ablation on Disentangled Inversion Technique
To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from the Table 3: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score. 2) The incorporation of target distance into target latent and harmonic branch leads to an improvement in structure and content integrity but the CLAP score sharply decreases. 3) Add the harmonic distance into the harmonic latent witness a distant decrease.
[…]
$[{\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ 14.28 0.17 2.64 4.67 0.57
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{src}]$ 13.17 0.17 2.55 4.59 0.58
$[{\bm{d}}_{src},{\bm{d}}_{tgt},{\bm{0}}]$ 11.24 0.15 2.44 4.32 0.56
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 4.29 0.56
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{har}]$ 37.51 0.27 27.2 14.23 0.28
$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.60
Table 3: Ablation results about disentangled inversion technique. [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance to the corresponding branch in Algorithm 2.
5.5 Qualitative Results
Figure 4: Visualizations of the source audio’s mel and edited mel-spectrograms by different editing methods.
To complement our quantitative findings, we present a qualitative comparison in Figure 4. Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with the preservation of melodic content and harmonic structure. In contrast, our Disentangled Inversion Control demonstrates superior performance in precise music editing while preserving structural integrity. We provide additional qualitative results in the Appendix G for further examination.
6 Conclusion
In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts. Our retrospective analysis revealed that while recent methods like DDIM inversion have propelled zero-shot editing forward, they are not without their shortcomings. We identified that DDIM inversion is plagued by an accumulation of errors throughout the diffusion steps, which detracts from its overall effectiveness. Moreover, the absence of a fine-tuned attention control mechanism was found to impede the nuanced editing of music. To address these issues, we introduced the Disentangled Inversion framework, a novel approach that meticulously separates the source and target branches to amplify their individual strengths in preservation and editing. In tandem, we presented the Harmonized Attention Control technique, which enhances the preservation of essential content by integrating the attention maps from the original audio into the diffusion process. Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,000 samples across 10 varied editing categories—attested to the superiority of our methods.
Appendix A Preliminaries And Analyses
This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied in diffusion models for text-guided audio synthesis. It further explores the challenges associated with these methods.
A.1 Diffusion Models
Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, which corresponds to the given conditioning prompt. We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
[…]
Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$. At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using pre-trained network for $T$ steps. To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM samling:
${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (11)
A.2 DDIM Inversion
While diffusion models have superior characteristics in the feature space that can support various down-stream tasks, it is hard to apply them to audios in the absence of natural diffusion feature space for non-generated audios. Thus, a simple inversion technique known as DDIM inversion is commonly used for unconditional diffusion models, predicated on the presumption that the ODE process can be reversed in the limit of infinitesimally small steps:
${\bm{z}}^{*}_{t}=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}{\bm{z}}^{*}_{t-1}+\sqrt{\alpha_{t}}(\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1})\epsilon_{\theta}({\bm{z}}^{*}_{t-1},t-1)$ (12)
However, in most text-based diffusion models, this presumption cannot be guaranteed, resulting in a perturbation from ${\bm{z}}_{t}$ to ${\bm{z}}^{*}_{t}$. Consequently, an additional perturbation from ${\bm{z}}^{*}_{t}$ to ${\bm{z}}^{src}_{t}$ arises when sampling an audio from ${\bm{z}}^{*}_{T}$ where $\alpha$ is hyper-parameter:
${\bm{z}}_{t}=\sqrt{\alpha_{t}}{\bm{z}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$ (13)
A.3 Classifier-free Guidance
Classifier-free Guidance (CFG) (Ho & Salimans, 2021) is proposed to overcome the limitation of text-conditioned models, where text adherence could be weak. However, a higher guidance scale, $\omega$, which is intended to strengthen the model’s fidelity to the text prompt, inadvertently magnifies the accumulated inversion error. This becomes problematic in editing scenarios where precise control over the audio synthesis is desired. The modified noise estimation in CFG can be expressed as:
$\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$ (14)
where $\varnothing$ = ("") is the embedding of a null text. This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{"}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
Appendix B Benchmark construction
B.1 General information
Here are details about our ZoME-Bench dataset (Zero-shot Music Editing Benchmark).This dataset contains 1,000 audio samples , which are selected from MusicCaps , and each is 10 seconds long with a sample rate of 16k.
We refactored their captions to express an edit and divide them into 10 parts , each representing a type of editing. Sample and details are shown in the following table
Editing type id Editing type size origin prompt editing prompt editing instruction
0 change instrument 131 ambient acoustic [guitar] music ambient acoustic [violin] music change the instrument from guitar to violin
1 add instrument 139 metal audio with a distortion guitar [and drums] metal audio with a distortion guitar add drums to the piece
2 delete instrument 133 an eerie tense instrumental featuring electronic drums [and synth keyboard] an eerie tense instrumental featuring electronic drums remove the synth keyboard
3 change genre 134 a recording of a solo electric guitar playing [blues] licks a recording of a solo electric guitar playing [rocks] licks change the genre from blues to rock
4 change mood 100 a recording featuring electric bass with an [upbeat] vibe a recording featuring electric bass with an [melancholic] vibe turn upbeat mood into melancholic mood
5 change rhythm 69 a live ukulele performance featuring [fast] strumming and emotional melodies a live ukulele performance featuring [slow] strumming and emotional melodies change fast rhythm into slow one
6 change background 95 female voices in unison with [acoustic] guitar female voices in unison with [electric] guitar switch acoustic guitar to electric guitar
7 change melody 121 this instrumental song features a [relaxing] melody this instrumental song features a [cheerful] melody change relaxing melody into cheerful melody
8 extract instrument 111 a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar] a reggae rhythm recording with bongos extract bongos from the recording
9 random 67 / / /
Table 4: Information of ZoME-Bench dataset
B.2 Annotation Process
We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023). With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type "change melody" as examples):
"There is a description of a Piece of music ,Please judge whether the description has information of melody.If not , just answer "Flase" ,else change its melody properly into the opposite one ,just change the adjective and don’t replace any instrument! "blended_word" is [origin melody, changed melody] "emphasize" is [changed melody] "blended_word" and "emphasize" are tuples.
Question:(A mellow, passionate melody from a noisy electric guitar) Answer:("original_prompt": "A mellow, [passionate] melody from a noisy electric guitar", "editing_prompt": "A mellow, [soft] melody from a noisy electric guitar", "blended_word": ["passionate melody","soft melody"], "emphasize" :["soft melody"]) Question:(A recording of solo harp music with a dreamy, relaxing melody.)
Answer:("original_prompt": "A recording of solo harp music with a dreamy, [relaxing] melody.", "editing_prompt": "A recording of solo harp music with a dreamy, [nervous] melody.", "blended_word": ["relaxing melody","nervous melody"], "emphasize" :["nervous melody"])
Question:("A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
Answer:("original_prompt": "A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", "editing_prompt": "A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", "blended_word": ["harmonized flute melody","passionate flute melody"]), "emphasize" :["passionate flute melody"])
now we have Question({origin caption})
Answer(?)"
In the same way , instructions were appended by prompt as follows(take type "change melody" as examples):
"There is two description of different pieces of music divided by &,Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
Answer: change relaxing melody into cheerful melody
Question:this song features acapella harmonies with a [high pitched] melody complemented by both high pitched female whistle tones and male low pitch tones & this song features acapella harmonies with a [smooth] melody complemented by both high pitched female whistle tones and male low pitch tones
Answer:turn a high pitched melody into smooth melody
Question: a traditional and hopeful song with a harmonizing throaty male vocal and [dissonant] background melody from strings albeit presented in low quality & a traditional and hopeful song with a harmonizing throaty male vocal and [harmonic] background melody from strings albeit presented in low quality
Answer: change dissonant melody into harmonic melody  Question: [’original prompt’] & [’editing prompt’]
Answer(?)"
Through this method accmopnied with rounds of manual review ,we guarantee the quality of this Benchmark.
B.3 Data Format
Taking the first piece as example , we express our data in json format with six keys
"$000$": {
"editing_prompt": "a live recording of ambient acoustic
[violin] music",
"original_prompt": "a live recording of ambient acoustic
[guitar] music",
"blended_word": "(\"guitar\", \"violin\")",
"emphasize": "(\"violin\")",
"audio_path": "wavs/MusicCaps_-4SYC2YgzL8.wav",
"editing_type_id": "0",
"editing_instruction": "change the instrument from guitar
to violin"
"Editing_prompt" is the edited caption while "Original_prompt" is the origin one . "Blended_word" is the subject to be edited."Emphasize" is the word which should be emphasized. "editing_instruction" is the description of editing.Something else that needs attention is in the editing type "delete instrument" , we create another key "neg_prompt" , which helps to reduce the possibility of deleted instruments still appearing.
Appendix C Implementation Details
For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023b). Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD. These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval. In line with the methodology described by  Manor & Michaeli (2024), we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion. For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12. The forward guidance of DIC is 1 while the reverse scale is 5. We chose these values by exploration about different guidance scale, as discussed in Appendix D. We conduct all experiments in NVIDIA 4090. For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023b). To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD, and conduct on an NVIDIA 4090. The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
Our methodology is aligned with the protocol established by  Manor & Michaeli (2024), where we have adopted a forward guidance scale of 3 and a reverse guidance scale of 12 for DDPM inversion. In contrast, the DDIM inversion employs a guidance scale of 5, and SDEdit utilizes a guidance scale of 12. For Disentangled Inversion Control, we have determined the forward guidance to be 1 and the reverse scale to be 5. These specific guidance scale values are selected after extensive experimental exploration, the details of which are discussed in Appendix D.
C.1 Metrics
There are details about five metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio. (2) CLAP Score (Elizalde et al., ): This criterion evaluates the degree to which the output conforms to the specified target prompt. (3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the original audio. (4) MSE (Mean Squared Error): This distance metric quantifies the variance between the source and the edited audio. (5) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
Appendix D Quantitative Results
For the lack of systematic experiments which determines the optimal combination of guidance scales for achieving the best editing performance, and analysis on how these guidance scales affect final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
Guidance Scale Structure Background Preservation CLIP Similariy
[…]
1 1 8.56 0.12 1.17 3.25 0.51
1 2.5 11.97 0.15 2.49 4.54 0.56
1 5 15.99 0.17 4.22 6.07 0.60
1 7.5 15.99 0.17 4.22 6.07 0.59
2.5 1 22.80 0.20 6.39 8.65 0.30
2.5 2.5 14.24 0.16 2.50 5.40 0.46
2.5 5 14.46 0.16 3.31 5.49 0.53
2.5 7.5 15.51 0.17 3.94 5.89 0.53
5 1 29.94 0.24 9.81 11.36 0.20
5 2.5 29.16 0.24 9.11 11.07 0.22
5 5 22.15 0.20 5.59 8.40 0.36
5 7.5 17.57 0.18 5.57 6.67 0.48
7.5 1 31.41 0.25 10.62 11.92 0.20
7.5 2.5 31.05 0.25 10.14 11.78 0.20
7.5 5 29.20 0.24 9.32 11.08 0.24
7.5 7.5 24.16 0.22 7.33 9.17 0.34
Table 5: Ablation Studies on Different Guidance Scale
Appendix E Potential Negative Societal Impacts
MEDIC may also lead to potential negative societal impacts that are worthy of consideration. If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MEDIC could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation.Hopefully , all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
Appendix F Limitations
In spite of remarkable outcome of our method , but due to the limitation of the generation model we used , we are incapable of instigating a profound change.
Though our method is faster than est-optimization and end-to-end diffusion models,but due to the numerous steps it requires (T=200), the duration of computing distance is quite long.Thus , we will implement a more powerful text-to-music generation model to support better editing, while trying to use consistency model or flow matching model to achieve high quality and fast music generation in future work
Appendix G Qualitative Results
For each type in ZoME-Bench , We provide samples to observe the capability of MEDIC intuitively.
G.1 Change Instrument
In the following figures , we show the capability of MEDIC to change the instrument.Here we edit the ground truth music piece with original prompt "a live recording of ambient acoustic [guitar] music" and editing prompt "a live recording of ambient acoustic [violin] music". The difference of instruments can be observed in the Mel-spectrum
source: a live recording of ambient acoustic [guitar] music
edit: a live recording of ambient acoustic [violin] music
Figure 5: Editing Type 0 :Change Instrument
G.2 Add Instrument
In the following figures , we show the capability of MEDIC to add more instrument.Here we edit the ground truth music piece with original prompt "a heavy metal instructional audio with a distortion guitar" and editing prompt "a heavy metal instructional audio with a distortion guitar [and drums]". The appearance of new instrument can be observed in the Mel-spectrum which present drum sound of high frequency.
source: a heavy metal instructional audio with a distortion guitar
edit: a heavy metal instructional audio with a distortion guitar [and drums]
Figure 6: Editing Type 1 :Add Instrument
G.3 Delete Instrument
In the following figures , we show the capability of MEDIC to delete instrument.Here we edit the ground truth music piece with original prompt "a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and editing prompt "a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood". The vanishing of instrument can be observed in the Mel-spectrum
source: a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood
edit: a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood
Figure 7: Editing Type 2 :Delete Instrument
G.4 Change Genre
In the following figures , we show the capability of MEDIC to change genre of a music piece. Here we edit the ground truth music piece with original prompt "a recording of a solo electric guitar playing [blues] licks" and editing prompt "a recording of a solo electric guitar playing [rock] licks". The obvious difference of genre can be observed in the Mel-spectrum
source: a recording of a solo electric guitar playing [blues] licks
edit: a recording of a solo electric guitar playing [rock] licks
Figure 8: Editing Type 3 :Change Genre
G.5 Change Mood
Mood is an important attribute of a music. In the following figures ,we show the capability of MEDIC to change mood of a music piece. Here we edit the ground truth music piece with original prompt "a recording of [aggressive] electronic and video game music with synthesizer arrangements"and editing prompt "a recording of [peaceful] electronic and video game music with synthesizer arrangements". The change of Mood can be observed in the Mel-spectrum
source: a recording of [aggressive] electronic and video game music with synthesizer arrangements
edit: a recording of [peaceful] electronic and video game music with synthesizer arrangements
Figure 9: Editing Type 4 :Change Mood
G.6 Change Rhythm
Rhythm represents the speed of the music. In the following figures ,we show the capability of MEDIC to change Rhythm of a music piece. Here we edit the ground truth music piece with original prompt "a [slow] tempo ukelele tuning recording with static"and editing prompt "a [fast] tempo ukelele tuning recording with static". The change of Rhythm can be observed in the Mel-spectrum.The edited Mel-spectrum is much intensive.
source: a [slow] tempo ukelele tuning recording with static
edit: a [fast] tempo ukelele tuning recording with static
Figure 10: Editing Type 5 :Change Rhythm
G.7 Change Background
In the following figures ,we show the capability of MEDIC to change background of instrument of a music piece. Here we edit the ground truth music piece with original prompt "an amateur ukulele recording with a [medium to uptempo] pace"and editing prompt "an amateur ukulele recording with a [steady and rhythmic] pace". The change of instrument background can be observed in the Mel-spectrum.
source: an amateur ukulele recording with a [medium to uptempo] pace"and editing prompt
edit: an amateur ukulele recording with a [steady and rhythmic] pace
Figure 11: Editing Type 6 :Change Background
G.8 Change Melody
In the following figures ,we show the capability of MEDIC to change melody of a music piece. Here we edit the ground truth music piece with original prompt "this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo"and editing prompt "this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo". The change of Melody can be observed in the Mel-spectrum
source: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
Figure 12: Editing Type 7 :Change Melody
G.9 Extract Instrument
In the following figures ,we show the capability of MEDIC to extract one certain instrument of a music piece. Here we edit the ground truth music piece with original prompt "a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]"and editing prompt "a reggae rhythm recording with bongos". The change of instruments can be observed in the Mel-spectrum. Much sounds vanished.
source: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
Figure 13: Editing Type 8 :Extract Instrument
Appendix H safeguards
In the processing of the data and models involved in this study, we fully considered the potential risks. We ensure that all data sources are rigorously screened and vetted,and the model we used is absolutely trained from the safe dataset to minimize security risks of being misused.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We accurately reflect the paper’s contributions and scope in the abstract and section 1 introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in section F limitations
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide theory assumptions and proofs in section 2 preliminaries and analyses
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: we will open-source our codes(contains model or their access method) and the benchmark we constructed, which is enough to reproduce our experiments.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: we specify all the training and test details in section 5 Experiments and section C Implementation Details
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We only use one random seed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide sufficient information on the computer resources in section 5.1 Implementation Details
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Positive impacts is in section 6 Conclusion.Negative impacts in section E Potential Negative Societal Impacts
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We take actions for safeguards , which is described in section H Safeguards
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Creators or original owners of assets, used in the paper, are properly credited and the license and terms of use explicitly mentioned and are properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our assets is well documented and the document is provided
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work do not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work do not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 We ensure that all data sources are rigorously screened and vetted,and the model we used is absolutely trained from the safe dataset to minimize security risks of being misused.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We accurately reflect the paper’s contributions and scope in the abstract and section 1 introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss the limitations in section F limitations
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: We provide theory assumptions and proofs in section 2 preliminaries and analyses
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: we will open-source our codes(contains model or their access method) and the benchmark we constructed, which is enough to reproduce our experiments.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: we specify all the training and test details in section 5 Experiments and section C Implementation Details
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34. Justification: We only use one random seed.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We provide sufficient information on the computer resources in section 5.1 Implementation Details
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Positive impacts is in section 6 Conclusion.Negative impacts in section E Potential Negative Societal Impacts
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54. Justification: We take actions for safeguards , which is described in section H Safeguards
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Creators or original owners of assets, used in the paper, are properly credited and the license and terms of use explicitly mentioned and are properly respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: Our assets is well documented and the document is provided
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: Our work do not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: Our work do not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
 Abstract
-Text-guided diffusion models catalyze a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts.
-Recent advancements introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pre-trained diffusion models for audio modification.
+Text-guided diffusion models make a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts.
+Recent works introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pretrained diffusion models for audio modification.
 Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy.
-And the lack of attention control hinders the fine-grained manipulations of music.
-To counteract these limitations, we introduce the Disentangled Inversion technique, which is designed to disentangle the diffusion process into triple branches, thereby magnifying their individual capabilities for both precise editing and preservation.
-Furthermore, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention and cross-attention with an additional Harmonic Branch to achieve the desired composition and structural information in the target music.
-Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing whilst safeguarding structural integrity.
-To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across 10 distinct editing categories, which facilitates both zero-shot and instruction-based music editing tasks.
-Our method demonstrates unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques.
-[^1] Audio samples are available at https://MEDIC-Zero.github.io/.
-[^2] Benchmark data are available at https://huggingface.co/datasets/liuhuadai/ZoME-Bench.
+Moreover, existing editing methods fail to achieve effective complex non-rigid music editing while maintaining essential content preservation and high editing fidelity.
+To counteract these issues, we introduce the Disentangled Inversion technique to disentangle the diffusion process into triple branches, rectifying the deviated path of the source branch caused by DDIM inversion.
+In addition, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention control and cross-attention control with an intermediate Harmonic Branch to progressively achieve the desired harmonic and melodic information in the target music.
+Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing while safeguarding content integrity.
+To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across ten distinct editing categories.
+This facilitates both zero-shot and instruction-based music editing tasks.
+Our method achieves unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques.
+[^1] The project webpage is available at https://MEDIC-Edit.github.io/.
 1 Introduction
-Advancements in text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have markedly enhanced audio generation, leveraging their impressive capability for realistic and varied outputs.
-These models (Liu et al., 2023b; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes in response to specific textual prompts.
-Early audio editing strategies rely on training models from scratch or test-time optimization, both of which are hampered by intensive computational demands.
-Recent advances (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through DDIM (Song et al., 2020) and DDPM (Ho et al., ) inversion techniques, but challenges remain.
-Key among these challenges is maintaining editing fidelity - ensuring that the model’s output matches the given instructions - and preserving essential content, a hallmark of effective audio editing that targets only necessary changes.
-Balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving suboptimal for conditional diffusions (Mokady et al., 2023).
-Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space, yet at the expense of reduced modification capabilities due to noise reduction.
+Text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have made great progress in audio generation (Evans et al., 2024a, b), leveraging their impressive capability for realistic and varied outputs.
+In particular, these models (Liu et al., 2023c; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes for specific textual prompts.
+Early audio editing strategies rely on training models from scratch (Copet et al., 2023; Agostinelli et al., 2023) or test-time optimization (Paissan et al., 2023; Plitsis et al., 2024), hampered by intensive computational demands.
+Recent works (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) and Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) inversion techniques, but challenges remain.
+Key among these challenges are maintaining fidelity of editing - ensuring the editing aligns with the provided instructions - and essential content preservation, ensuring that the unaltered musical attributes in the target prompts remain unchanged.
+However, balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving sub-optimal for conditional diffusion models (Mokady et al., 2023).
+Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space.
+However, this comes at the expense of reduced modification capabilities due to noise reduction.
 In this work, we rigorously examine the shortcomings of the DDIM inversion approach.
-Our comprehensive analysis, suggests that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence.
-The implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) further amplifies the accumulated errors.
-Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has demonstrated its power in achieving high fidelity and essential content preservation.
-Yet, analogous control mechanisms in music editing are noticeably absent.
-To bridge this gap, we introduce an innovative Disentangled Inversion Control technique.
-This technique is comprised of two main components: Harmonic Attention Control and Disentangled Inversion.
-The former integrates cross-attention and mutual self-attention controls, supplemented by a novel Harmonic Branch, which facilitates precise, fine-grained manipulations for both rigid and non-rigid editing scenarios.
-The latter component strategically deconstructs the diffusion process into three distinct branches, enabling each to realize its full potential independently.
-Acknowledging the lack of standardized benchmarks in the field of audio editing, we introduce ZoME-Bench.
-The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories.
+Our comprehensive analysis indicates that while techniques like DDIM inversion provide an editable foundation for audio synthesis, they lack precision and may compromise the integrity of the original audio.
+The primary issue stems from the assumption of perfect reversibility in the ordinary differential equation (ODE) process, which is frequently not met during text-conditional editing.
+Consequently, this leads to distortions during the inversion.
+And the implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) aims to improve text adherence.
+However, it inadvertently amplifies the accumulated errors from the inversion process.
+Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has shown promise in achieving high fidelity and essential content preservation.
+For instance, MusicMagus (Zhang et al., 2024) introduces Cross-Attention Control for fine-grained music manipulation of rigid tasks.
+Nevertheless, these methods fail to resolve the issues of accumulated errors and struggle to achieve accurate editing for both rigid and non-rigid tasks, as illustrated in Figure 1.
+We introduce an innovative Disentangled Inversion Control technique to bridge this gap.
+This technique has two main components: Harmonic Attention Control and Disentangled Inversion.
+Cross-attention control (Hertz et al., 2022) and mutual self-attention control (Cao et al., 2023) have demonstrated robust editing capabilities for rigid and non-rigid image editing tasks, respectively.
+However, blindly combining these two approaches sequentially for music editing can result in sub-optimal outcomes, particularly in the original dual-branch setup, where it struggles with global attention refinement.
+To address this challenge, we introduce an intermediate branch called the Harmonic Branch, designed to modify both rigid and non-rigid attributes in music progressively.
+Furthermore, we disentangle the diffusion process into triple branches, correcting the deviation path caused by CFG in the source branch, which affects the essential content preservation.
+The other branches remain unchanged to ensure the highest possible edit fidelity.
+Due to a lack of standardized benchmarks in audio editing, we introduce ZoME-Bench.
+The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories across rigid and non-rigid tasks.
 Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing.
-Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
+Experimental results on ZoME-Bench indicate that MEDIC outperforms baselines, achieving significant improvements in essential content preservation and editing fidelity.
+Additionally, MEDIC demonstrates state-of-the-art performance in the variable-length music editing settings of the MusicDelta dataset.
 Our contributions can be summarized as follows.
-1) We introduce a novel, tuning-free methodology referred to as Disentangled Inversion Control (DIC), designed to facilitate both consistent manipulations of musical elements and intricate non-rigid editing tasks.
-2) An Harmonic Attention Control framework is introduced to unify cross-attention control and self-attention control, which enables both rigid and non-rigid editing.
-3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error without optimization by branch disentanglement, aiding in accurately editing the music while preserving the structural information.
-4) We introduce a tailored benchmark for music editing, named ZoME-Bench, and Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
+1) We introduce a novel, training-free methodology called Disentangled Inversion Control (DIC), designed to facilitate consistent manipulations of musical elements and intricate non-rigid editing tasks.
+2) An Harmonized Attention Control framework is introduced to unify cross-attention and mutual self-attention control, which enables both rigid and non-rigid editing.
+3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error by branch disentanglement and correction, aiding in accurately editing the music while preserving the content information.
+4) We build a new benchmark for music editing, named ZoME-Bench, which supports both zero-shot and instruction-based music editing.
 2 Related Works
-2.1 Text-based Audio editing
+Text-based Audio editing.
 The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt.
-Two primary challenges command intense focus.
-The first is the fidelity of editing, which pertains to the model’s ability to strictly follow the initial instructions.
-The second challenge involves essential content preservation, a critical attribute of competent audio editing models that ensures only suitable regional modifications are made during the editing process.
 Existing methodologies for addressing these intricate challenges typically follow one of three paths.
@@
 Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load.
-The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020) inversion strategies to extract diffusion noise vectors that match the source signal.
+The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020; Zhang et al., 2024) inversion strategies to extract diffusion noise vectors that match the source signal.
 Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework.
-In this study, we propose a new inversion technique named Disentangled Inversion Control.
+In this work, we propose a new inversion technique named Disentangled Inversion Control.
 This technique aims to achieve accurate editing while preserving structural information.
-2.2 Inversion Techniques
+Inversion Techniques.
 The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021).
@@
 A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges.
-For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during the editing process.
+For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during editing.
 Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024).
-Recently, optimization-based inversion methods using specific latent variables have gained popularity (Ju et al., 2023; Kawar et al., 2023).
+Optimization-based inversion methods using specific latent variables have recently gained popularity (Ju et al., 2023; Kawar et al., 2023).
 These are designed to minimize accumulated errors stemming from the DDIM inversion.
@@
 Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
-Figure 1: The framework and rationale behind Disentangled Inversion Control.
-Traditional DDIM inversion introduces incremental errors at each step, leading to only a coarse approximation of the source audio’s distribution.
-The implementation of classifier-free guidance exacerbates the error accumulation.
-In contrast, our method intricately separates the diffusion process into a triple-branch system, allowing each branch to fully realize its individual potential.
+Figure 1: Comparisons of our method with two-branch inversion techniques, including DDIM Inversion and MusicMagus.
+(a) Framework of DDIM Inversion, showing configurations with and without classifier-free guidance.
+(b) Framework of MusicMagus, which incorporates cross-attention control.
+(c) Framework of our method, featuring disentangled inversion control.
+(d) An illustration comparing the output of the two-branch techniques with the progressive output of our triple-branch approach.
 3 Disentangled Inversion Control
+3.1 Problem Definition and BenchMark Construction
+Despite significant work in text-to-audio generation models, particularly with the emergence of latent diffusion models (LDM), research on zero-shot music editing remains limited.
+Zero-shot music editing seeks to leverage the capabilities of text-to-music generation models to synthesize the desired music, denoted as ${\bm{x}}^{tgt}_{0}$.
+This synthesized music should align with the target edited text prompt ${\mathcal{P}}^{*}$, which is directly modified from the source music ${\bm{x}}^{src}_{0}$ and its corresponding text prompt ${\mathcal{P}}$.
+We compress source audio signals into latent ${\bm{z}}^{src}_{0}$ for inversion.
+To systematically validate our proposed method as a plug-and-play strategy for editing models, compare our method with existing zero-shot music editing methods, and compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset named ZoME-Bench (Zero-shot Music Editing Benchmark).
+ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps (Doh et al., 2023), spanning ten different editing types that include both rigid and non-rigid modifications.
+Each sample is accompanied by its corresponding source prompt, target prompt, human instruction, and source audio.
+Additionally, we include annotations relevant to attention control, such as blended words, to facilitate detailed evaluations. Further details about our benchmark can be found in Appendix B.
 3.2 Motivation
-Figure 1 and Preliminaries in Appendix 2 reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence.
+Figure 1 and Preliminaries in Appendix A reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision.
+This potentially compromising the essential content preservation.
 The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
@@
 (2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$.
-This sparks a plethora of attention control techniques that direct the image domain’s target denoiser network $\hat{\epsilon}_{\theta}$ to better reflect target prompts.
-Yet, analogous control mechanisms in music editing are noticeably absent.
-Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach poised to achieve fine-grained zero-shot music editing.
-DIC strategically disentangles the diffusion process as a triple-branch, granting each branch to optimize its functionality.
+This leads to the development of various attention control techniques that guide the target denoiser network $\hat{\epsilon}_{\theta}$ in the image domain to better align with target prompts.
+Yet, analogous control mechanisms for non-rigid music editing are noticeably limited.
+Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach to achieve both rigid and non-rigid music editing.
+DIC strategically disentangles the diffusion process as triple branches, allowing each branch to optimize its functionality.
 At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance.
@@
 3.3 Harmonized Attention Control Framework
-The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023a).
+The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023b).
 At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer.
@@
 $\text{Attention}(Q,K,V)=MV=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (1)
-where $Q$ denotes the query features derived from the audio features, $K$, $V$ represent the key and value features, and $M$ is the attention maps.
-We explore varying semantic transformations of audio content through bifurcated attention control strategies—cross-attention for rigid changes and mutual self-attention for non-rigid adjustments.
-Finally, we introduce an intermediate branch to host the desired composition and structural information in the target music.
+where $Q$ denotes the query features derived from the audio features, $K$ and $V$ represent the key and value features, and $M$ is the attention map.
+We explore varying semantic transformations of audio content through harmonized attention control strategies — cross-attention control for rigid changes and mutual self-attention control for non-rigid adjustments.
+Finally, we introduce an intermediate branch to host the desired harmonic and melodic information in the target music.
+The framework of harmonized attention control is depicted in Figure 2.
 Figure 2: The framework of Harmonized Attention Control (HAC).
@@
 3.3.1 Cross-Attention Control
-Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the modified prompt ${\mathcal{P}}^{*}$.
-We follow the practice of Prompt-to-Prompt and define CAC as Global Attention Refinement and Local Attention Blend.
+Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the target prompt ${\mathcal{P}}^{*}$.
+We follow the practice of Prompt-to-Prompt (Hertz et al., 2022) and define CAC as Global Attention Refinement and Local Attention Blend.
 Global Attention Refinement
@@
 $\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$ (6)
-For simplificity, we define the process of local editing in Equation 10 as:
+For simplicity, we define the process of local editing in Equation 5 as:
 ${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$ (7)
@@
 The approach is defined as:
-$\text{CrossEdit}(M_{t},M^{*}_{t},t)=\begin{cases}\text{Refine}(M^{src},M^{tgt})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
+$\text{CrossEdit}(M^{src},M^{tgt},t)=\begin{cases}\text{Refine}(M^{src}_{t},M^{tgt}_{t})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}_{t}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
 3.3.2 Mutual Self-Attention Control
@@
 3.3.3 Harmonic Branch Integretion
-The naive combination of cross-attention control and mutual self-attention control sequentially would lead to a sub-optimal outcome in the original dual-branch setup, especially failing the global attention refinement.
+The naive combination of cross-attention control and mutual self-attention control sequentially would lead to sub-optimal results in the original dual-branch setup, especially failing the global attention refinement.
 To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music.
-Our unified framework is detailed in Algorithm 2 and illustrated in Figure 2.
+Our unified framework is detailed in Algorithm 2.
 During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$.
-Following this, cross-attention control is applied on $M_{har}$ and $M_{tgt}$ to refine the semantic information for $M_{tgt}$.
+This lays the formal structure of the target music.
+Following this, cross-attention control is applied on $M^{har}$ and $M^{tgt}$ to refine the semantic information for $M^{tgt}$.
+As illustrated in Figure 2, the harmonic branch output $z^{har}_{0}$ reflects the requested non-rigid changes (e.g., “violin"), while preserving the rigid content semantics (e.g., “with noise").
+The target branch output $z^{tgt}_{0}$ builds upon the structural layout of the $z^{har}$ while reflecting the requested rigid changes (e.g., “with noise").
 Algorithm 1 Harmonized Attention Control in one DDIM Forward Process
@@
 Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal.
-Increasing the guidance scale enhances editability but sacrifices reconstruction accuracy due to latent code deviation during editing.
-In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the target, and the harmonized branch.
-This bifurcation is designed to optimize the performance of each branch separately.
+Increasing the guidance scale enhances editability, but sacrifices reconstruction accuracy due to latent code deviation during editing.
+In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the harmonic, and the target branch.
+This decoupling is designed to unleash the capabilities of each branch separately.
 For the source branch, we implement a targeted correction mechanism.
@@
 This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio.
-On the other hand, the target branch and harmonic branch is left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio.
+On the other hand, the target branch and harmonic branch are left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio.
 The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio.
@@
 11:  return  ${\bm{z}}^{tgt}_{0}$
-4 BenchMark Construction
-Despite recent significant advancements in music generation utilizing the latent diffusion model, music editing has received comparably less attention, especially zero-shot music editing.
-Hence, to systematically validate our proposed method as a plug-and-play strategy for editing models and compare our method with existing zero-shot music editing methods, as well as compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset, named ZoME-Bench (Zero-shot Music Editing Benchmark).
-ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps, spanning ten distinct editing types such as add instrument, delete instrument, change genre, and so on.
-Each sample in this collection is accompanied by its corresponding source prompt, target prompt, human instruction, and the source audio.
-Additionally, annotations pertinent to attention control methodologies, such as blended words, are meticulously included to facilitate detailed evaluations.
-We attach the details about our benchmark in Appendix B.
 4 Experiments
-5.1 Implementation Details
-We perform the inference of different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023b) models with 200 inference steps.
-The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA 4090Ti GPU for a fair comparison.
-We evaluate all methods in ZoMo-Bench across all editing types.
+Implementation Details.
+We infer different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023c) models with 200 inference steps.
+The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA A800 GPU for a fair comparison.
+We evaluate all methods in ZoMo-Bench across all editing types with fixed length.
+And we use the MusicDelta subset of the MelodyDB (Bittner et al., 2014) dataset for variable length comparisons, comprised of 34 musical excerpts in varying styles and in lengths ranging from 20 seconds to 5 minutes.
 Details can be found in Appendix C.
-5.2 Evaluation Metrics
-In order to evaluate the performance of our novel Disentangled Inversion Control framework, we employ a diverse set of five metrics that encompass three critical dimensions of audio quality and fidelity: Struture Distance,CLAP Score,LPAPS,MSE (Mean Squared Error)and FAD (Fréchet Audio Distance).
-All the metrics are introduced in detail in Appendix C.1.
+Evaluation Metrics.
+Our models employ a comprehensive evaluation using both objective and subjective metrics to assess essential content preservation, text-audio alignment fidelity, and audio quality.
+Objective metrics include Structure Distance (SD) (Ju et al., 2023), CLAP (Contrastive Language-Audio Pretraining) Score (Elizalde et al., 2023), LPAPS (Learned Perceptual Audio Patch Similarity) (Iashin & Rahtu, 2021; Paissan et al., 2023), and FAD (Fréchet Audio Distance) (Kilgour et al., 2018).
+The CLAP Score evaluates how well the edited audio aligns with the target prompt, while Structure Distance, LPAPS, and FAD, which have been adapted from image domain metrics, measure the similarity between the edited and source audio.
+Specifically, LPAPS quantifies the consistency of the edited audio in relation to the source audio, and the FAD metric assesses the distance between two distributions of audio signals.
+The Structure Distance measures the structural similarity between the edited and source audio.
+For the subjective evaluation, we conduct crowd-sourced human assessments using the Mean Opinion Score (MOS) to evaluate both editing fidelity (MOS-Q) and content preservation (MOS-P).
+We attach the details of all metrics in Appendix C.1.
 4.1 Zero-shot Music Editing Results
+Method Objective Metrics Subjective Metrics
 […]
-SDEdit 25.87 0.22 12.18 9.82 0.40
-DDPM-Friendly 22.36 0.21 14.08 8.49 0.41
-DDIM Inversion 17.11 0.18 6.24 6.12 0.55
-DIC 11.97 0.15 2.49 4.54 0.60
-Table 1: Comparison with baselines on ZoMo-Bench about struture and content preservation and CLAP similiarity. DIC: our proposed Disentangled Inversion Control.
-We present a comparative study of our Disentangled Inversion Control(DIC) against several established music editing techniques, namely SDEdit (Liu et al., ), DDIM Inversion Ho et al.
-, and DDPM-Friendly (Manor & Michaeli, 2024).
-Utilizing the ZoMo-Bench test set we released, we assess the quality of the generated audio samples.
-The assessment primarily concentrates on two key aspects: Essential Content Preservation and Editing Fidelity.
-The results are summarized in Table 1, leading to two main conclusions:
-(1) Our Disentangled Inversion Control demonstrates substantial advancements in maintaining essential content and enhancing editing fidelity, in comparison to the baseline methods.
-(2) Despite DDIM Inversion achieving high CLAP similarity scores, significantly underperforms in preserving essential content when compared to our approach.
-(a) Spider chart for editing quality.
-(b) Spider chart for editing consistency.
+AudioLDM 2 23.86 0.21 10.36 0.58 73.48 70.12
+MusicGen 23.39 0.21 6.63 0.59 75.46 71.28
+SDEdit 25.87 0.22 12.18 0.40 69.38 66.23
+DDIM Inversion 22.52 0.21 9.51 0.49 73.10 3.38
+MusicMagus 16.23 0.19 5.15 0.55 75.12 74.34
+DDPM-Friendly 18.30 0.19 5.16 0.54 75.27 73.86
+MEDIC 11.97 0.15 2.49 0.61 79.81 77.29
+Table 1: Comparison with baselines on ZoME-Bench with fixed length about structure, content preservation, and CLAP similarity.
+We present a comparative study of our Disentangled Inversion Control (DIC) against several established music generation and editing baselines, including AudioLDM 2 (Liu et al., 2023c), MusicGen (Copet et al., 2023), SDEdit (Liu et al., 2023a), DDIM Inversion Ho et al.
+(2020), MusicMagus (Zhang et al., 2024), and DDPM-Friendly (Manor & Michaeli, 2024).
+Utilizing the ZoMo-Bench test set we developed, we assess the quality of the generated audio samples, focusing on two key aspects: Content Preservation and Editing Fidelity.
+The results, summarized in Table 1, lead to the following conclusions: (1) Our method, MEDIC, significantly outperforms both generation-based and inversion-based models in terms of editing fidelity and audio similarity, demonstrating its effectiveness in addressing complex editing tasks.
+(2) Although DDPM-Friendly and MusicMagus show improvements in essential content preservation, they struggle to maintain high text-audio alignment compared to generation models.
+MEDIC effectively addresses accumulated errors, achieving high audio similarity and an impressive CLAP score, reflecting its superior editing fidelity.
+(3) MEDIC also excels in subjective metrics, achieving the best performance in audio quality and content preservation.
+Variable Length Comparisons We further evaluate MEDIC against baseline methods in a variable length setting, with the results presented in Table 2.
+Our analysis leads to the following conclusions: (1) MEDIC outperforms all baselines across all metrics, demonstrating the effectiveness of our approach in variable-length scenarios.
+(2) Inversion-based baselines experience a significant degradation in CLAP score, while MEDIC maintains the highest CLAP score of 0.51 on the MusicDelta dataset.
+Method Objective Metrics Subjective Metrics
+[…]
+AudioLDM 2 24.40 0.22 7.07 0.44 66.37 64.28
+MusicGen 27.71 0.23 7.70 0.46 67.41 63.76
+SDEdit 28.12 0.24 13.21 0.24 62.80 62.18
+DDIM Inversion 23.5 0.21 10.12 0.27 65.94 65.73
+MusicMagus 25.6 0.22 7.13 0.43 67.45 67.12
+DDPM-Friendly 21.53 0.23 6.68 0.30 66.34 67.28
+MEDIC 19.5 0.20 6.58 0.51 71.62 70.18
+Table 2: Comparison with baselines in variable length setting.
+(a) Spider chart for CLAP score comparisons .
+(b) Spider chart for FAD score comparisons.
 Figure 3: A comprehensive performance evaluation on the ZoME-bench.
-We present spider charts of editing quality (CLAP Scores) and consistency (LPAPS) across 9 editing tasks for DDPM-Friendly, DDIM Inversion, and Disentangled Inversion Control methods.
-To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing.
-The LPAPS and CLAP scores are depicted in Figure 3.
+We present spider charts of CLAP scores and FAD scores across 10 editing tasks for DDPM-Friendly, DDIM Inversion, MusicMagus, and Disentangled Inversion Control methods.
+Fine-grained Comparisons on ZoME-Bench To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing.
+The FAD and CLAP scores are depicted in Figure 3.
 The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks.
-(2) While methods like DDIM Inversion exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing fine-grained manipulations, as reflected in their inferior performance, especially in the "Delete Instrument" and "Change Melody" non-rigid editing type.
+(2) While baselines exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing non-rigid manipulations, as reflected in their inferior performance, especially in the “Change Genre" and “Change Melody" non-rigid editing types.
 4.2 Ablation Study
-The ablation studies presented in this section aim to validate the contributions of the Unified Attention Control and Refining Dual-path Inversion to our framework’s overall performance.
-Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in the Appendix D.
+The ablation studies presented in this section aim to validate the contributions of the Harmonized Attention Control and Disentangled Inversion to our framework’s overall performance.
+Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in Appendix D.
 4.2.1 Ablation on Attention Control Methods
 […]
-HAC 11.97 0.15 2.49 4.54 0.60
-w/ CAC 10.24 0.19 2.75 4.76 0.56
-w/ MSAC 9.92 0.18 2.50 4.66 0.58
-w/ HB 11.75 0.16 2.59 4.67 0.59
-Table 2: Ablation results about different control methods. HAC: Harmonized Attention Control.
-To ascertain the impact of our attention control mechanisms, we perform ablations on the following configurations: Cross-attention Control Only (w/ CAC), Mutual Self-attention Control only (w/ MSAC), without Harmonic Branch (w/o HB), and any attention control (w/o Control).
+HAC 11.97 0.15 2.49 0.61
+w/o MSA Control 14.13 0.19 2.75 0.56
+w/o CA Control 13.78 0.18 2.50 0.58
+w/o Harmonic Branch 12.75 0.16 2.59 0.59
+Table 3: Ablation results about different control methods. MSA: Mutual Self-Attention, CA: Cross Attention.
+To validate the impact of our attention control mechanisms, we perform ablations on the following configurations: Remove Mutual Self Attention Control (w/o MSA Control), Remove Cross Attention Control (w/o CA Control), and without Harmonic Branch (w/o Harmonic Branch).
 The findings, detailed in Table 3, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance.
-(2) The combined use of CAC and MSAC further improves results, but the absence of the harmonic branch yields suboptimal outcomes.
+(2) Although the naive combination of mutual self-attention control and cross-attention control improves the preservation and fidelity, it still yields sub-optimal outcomes due to the lack of a progressive harmonic branch.
+This demonstrates the effectiveness of harmonized attention control.
 4.2.2 Ablation on Disentangled Inversion Technique
-To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from the Table 3: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score.
-2) The incorporation of target distance into target latent and harmonic branch leads to an improvement in structure and content integrity but the CLAP score sharply decreases.
-3) Add the harmonic distance into the harmonic latent witness a distant decrease.
+To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from Table 4: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score.
+2) Incorporating target distance into the target branch and harmonic branch improves structural integrity and content preservation.
+However, this modification leads to a significant decrease in the CLAP score, indicating that while structural quality may improve, the alignment with the target prompt suffers.
+3) Adding harmonic distance to the harmonic branch leads to a noticeable decline in performance.
+This finding showcases that introducing excessive deviations in the harmonic branch may adversely affect the overall audio quality and coherence.
 […]
@@
 $[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 4.29 0.56
-$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{har}]$ 37.51 0.27 27.2 14.23 0.28
-$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.60
-Table 3: Ablation results about disentangled inversion technique. [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance to the corresponding branch in Algorithm 2.
+$[{\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$ 37.51 0.27 27.2 14.23 0.28
+$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.61
+Table 4: Ablation results from the disentangled inversion technique.
+[,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance (line 6 in Algorithm 2).
+MSE means the mean square error loss between the edited audio features and source audio features.
 4.3 Qualitative Results
@@
 To complement our quantitative findings, we present a qualitative comparison in Figure 4.
-Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with the preservation of melodic content and harmonic structure.
-In contrast, our Disentangled Inversion Control demonstrates superior performance in precise music editing while preserving structural integrity.
+Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with preserving melodic content and harmonic structure.
+In contrast, our Disentangled Inversion Control performs better in precise music editing while preserving structural integrity.
 We provide additional qualitative results in the Appendix G for further examination.
@@
 In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts.
-Our retrospective analysis revealed that while recent methods like DDIM inversion have propelled zero-shot editing forward, they are not without their shortcomings.
-We identified that DDIM inversion is plagued by an accumulation of errors throughout the diffusion steps, which detracts from its overall effectiveness.
-Moreover, the absence of a fine-tuned attention control mechanism was found to impede the nuanced editing of music.
-To address these issues, we introduced the Disentangled Inversion framework, a novel approach that meticulously separates the source and target branches to amplify their individual strengths in preservation and editing.
-In tandem, we presented the Harmonized Attention Control technique, which enhances the preservation of essential content by integrating the attention maps from the original audio into the diffusion process.
-Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,000 samples across 10 varied editing categories—attested to the superiority of our methods.
+We proposed the Disentangled Inversion Control to support both rigid and non-rigid editing tasks.
+Instead of a two-branch setting, we add an intermediate branch named the harmonic branch to progressively integrate harmonic and melodic information in music by cross-attention control and mutual self-attention control.
+To counteract the accumulated errors caused by DDIM inversion and CFG, we introduced a simple but effective method named disentangled inversion to separate the diffusion process into triple branches and eliminate the latent discrepancy distance in the source branch.
+Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,100 samples across 10 varied editing categories—attested to the superiority of our methods.
+And the experiment results on variable length test sets and ablation studies further validated the effectiveness and robustness of our method.
+We envisage that our work could serve as a basis for future zero-shot music editing studies.
 Appendix A Preliminaries And Analyses
-This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied in diffusion models for text-guided audio synthesis.
+This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied to diffusion models for text-guided audio synthesis.
 It further explores the challenges associated with these methods.
 A.1 Diffusion Models
-Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, which corresponds to the given conditioning prompt.
+Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, corresponding to the given conditioning prompt.
 We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the Gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
@@
 Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$.
-At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using pre-trained network for $T$ steps.
-To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM samling:
+At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using a pre-trained network for $T$ steps.
+To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM sampling:
 ${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (11)
@@
 $\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$ (14)
-where $\varnothing$ = ("") is the embedding of a null text.
-This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{"}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
+where $\varnothing$ = (“”) is the embedding of a null text.
+This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{{}^{\prime\prime}}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
 Appendix B Benchmark construction
 B.1 General information
-Here are details about our ZoME-Bench dataset (Zero-shot Music Editing Benchmark).This dataset contains 1,000 audio samples , which are selected from MusicCaps , and each is 10 seconds long with a sample rate of 16k.
-We refactored their captions to express an edit and divide them into 10 parts , each representing a type of editing. Sample and details are shown in the following table
+Here are the details of our ZoME-Bench dataset (Zero-shot Music Editing Benchmark).
+This dataset contains 1,000 audio samples, selected from MusicCaps, with each sample being 10 seconds long and having a sample rate of 16k.
+We refactor the original captions to express specific edits and divide them into 10 parts, each representing a different type of editing. A sample and details are shown in the following table 5.
 Editing type id Editing type size origin prompt editing prompt editing instruction
@@
 We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023).
-With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type "change melody" as examples):
-"There is a description of a Piece of music ,Please judge whether the description has information of melody.If not , just answer "Flase" ,else change its melody properly into the opposite one ,just change the adjective and don’t replace any instrument!
-"blended_word" is [origin melody, changed melody] "emphasize" is [changed melody] "blended_word" and "emphasize" are tuples.
-Question:(A mellow, passionate melody from a noisy electric guitar) Answer:("original_prompt": "A mellow, [passionate] melody from a noisy electric guitar", "editing_prompt": "A mellow, [soft] melody from a noisy electric guitar", "blended_word": ["passionate melody","soft melody"], "emphasize" :["soft melody"]) Question:(A recording of solo harp music with a dreamy, relaxing melody.)
-Answer:("original_prompt": "A recording of solo harp music with a dreamy, [relaxing] melody.", "editing_prompt": "A recording of solo harp music with a dreamy, [nervous] melody.", "blended_word": ["relaxing melody","nervous melody"], "emphasize" :["nervous melody"])
-Question:("A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
-Answer:("original_prompt": "A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", "editing_prompt": "A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", "blended_word": ["harmonized flute melody","passionate flute melody"]), "emphasize" :["passionate flute melody"])
-now we have Question({origin caption})
-Answer(?)"
-In the same way , instructions were appended by prompt as follows(take type "change melody" as examples):
-"There is two description of different pieces of music divided by &,Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
+With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type “change melody" as examples):
+Description: “There is a description of a Piece of music, Please judge whether the description has information of melody.
+If not, just answer “Flase", else change its melody properly into the opposite one, just change the adjective and don’t replace any instrument!
+",“blended_word" is [origin melody, changed melody], “emphasize" is [changed melody], “blended_word" and “emphasize" are tuples.
+Question: (A mellow, passionate melody from a noisy electric guitar)
+Answer:(“source_prompt": “A mellow, [passionate] melody from a noisy electric guitar", “editing_prompt": “A mellow, [soft] melody from a noisy electric guitar", “blended_word": [“passionate melody", “soft melody"], “emphasize": [“soft melody"])
+Question: (A recording of solo harp music with a dreamy, relaxing melody.)
+Answer: (“source_prompt": “A recording of solo harp music with a dreamy, [relaxing] melody.", “editing_prompt": “A recording of solo harp music with a dreamy, [nervous] melody.", “blended_word": [“relaxing melody",“nervous melody"], “emphasize" :[“nervous melody"])
+Question: (“A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
+Answer: (“source_prompt": “A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", “editing_prompt": “A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", “blended_word": [“harmonized flute melody",“passionate flute melody"]), “emphasize" :[“passionate flute melody"])
+Now we have Question:({origin caption}), Answer(?)"
+In the same way, instructions are appended by prompt as follows (take type “change melody" as examples):
+Description: “There are two descriptions of different pieces of music divided by &, Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
 Answer: change relaxing melody into cheerful melody
@@
 Question: a traditional and hopeful song with a harmonizing throaty male vocal and [dissonant] background melody from strings albeit presented in low quality & a traditional and hopeful song with a harmonizing throaty male vocal and [harmonic] background melody from strings albeit presented in low quality
-Answer: change dissonant melody into harmonic melody  Question: [’original prompt’] & [’editing prompt’]
-Answer(?)"
-Through this method accmopnied with rounds of manual review ,we guarantee the quality of this Benchmark.
+Answer: change dissonant melody into harmonic melody
+Now we have Question: [‘source prompt’] & [‘editing prompt’], Answer(?)"
+Through this method, supplemented by rounds of manual review, we ensure the quality of this benchmark.
 B.3 Data Format
-Taking the first piece as example , we express our data in json format with six keys
+Taking the first piece as an example, we express our data in JSON format with six keys
 "$000$": {
@@
 [violin] music",
-"original_prompt": "a live recording of ambient acoustic
+"source_prompt": "a live recording of ambient acoustic
 [guitar] music",
@@
 to violin"
-"Editing_prompt" is the edited caption while "Original_prompt" is the origin one .
-"Blended_word" is the subject to be edited."Emphasize" is the word which should be emphasized.
-"editing_instruction" is the description of editing.Something else that needs attention is in the editing type "delete instrument" , we create another key "neg_prompt" , which helps to reduce the possibility of deleted instruments still appearing.
+“Editing_prompt" refers to the edited caption, while “source_prompt" denotes the original caption.
+"Blended_word" indicates the subject to be edited, and “Emphasize" represents the word that should be highlighted.
+“Editing_instruction" provides a description of the editing process.
+Additionally, in the editing type “delete instrument," we introduce another key, “neg_prompt", which helps reduce the likelihood of deleted instruments reappearing.
 Appendix C Implementation Details
-For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023b).
+For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023c).
 Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD.
@@
 For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12.
-The forward guidance of DIC is 1 while the reverse scale is 5.
-We chose these values by exploration about different guidance scale, as discussed in Appendix D.
+The forward guidance of MEDIC is 1 while the reverse scale is 5.
+We chose these values by exploring different guidance scales, as discussed in Appendix D.
 We conduct all experiments in NVIDIA 4090.
-For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023b).
-To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD, and conduct on an NVIDIA 4090.
+For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023c).
+To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD and conduct on an NVIDIA 4090.
 The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
@@
 C.1 Metrics
-There are details about five metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio.
-(2) CLAP Score (Elizalde et al., ): This criterion evaluates the degree to which the output conforms to the specified target prompt.
-(3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the original audio.
-(4) MSE (Mean Squared Error): This distance metric quantifies the variance between the source and the edited audio.
+Objective Metrics There are details about four metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) CLAP Score (Elizalde et al., 2023): This criterion evaluates the degree to which the output conforms to the specified target prompt.
+(2) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio.
+(3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the source audio.
 (4) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
+Subjective Metrics To directly reflect the quality of the audio generated, we carry out MOS (Mean Opinion Score) tests.
+These tests involve scoring two aspects: MOS-Q, which assesses the edited quality of the audio, and MOS-P, which measures the content preservation of edited audio.
+For assessing editing fidelity, the evaluators were specifically directed to “Does the natural language description align with the audio?” They were provided with both the audio and its corresponding caption.
+They were then asked to give their subjective rating (MOS-Q) on a 20-100 Likert scale.
+To assess essential content preservation, human evaluators were presented with source audio, target audio, source prompt, and target prompt.
+They were then asked to answer the question, “To what extent does the target audio retain the essential features of the source audio, such as melody, instrumentation, and overall style?” The raters had to select one of the options: “completely,” “mostly,” or “somewhat,” using a 20-100 Likert scale for their response.
+Our crowd-sourced subjective evaluation tests were conducted via Amazon Mechanical Turk where participants were paid \$8 hourly.
 Appendix D Quantitative Results
-For the lack of systematic experiments which determines the optimal combination of guidance scales for achieving the best editing performance, and analysis on how these guidance scales affect final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
+Analyses on Different CFG Scale The lack of systematic experiments that determine the optimal combination of guidance scales for achieving the best editing performance, and analysis of how these guidance scales affect the final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
 Guidance Scale Structure Background Preservation CLIP Similariy
@@
 1 2.5 11.97 0.15 2.49 4.54 0.56
-1 5 15.99 0.17 4.22 6.07 0.60
+1 5 15.99 0.17 4.22 6.07 0.61
 1 7.5 15.99 0.17 4.22 6.07 0.59
@@
 Table 6: Ablation Studies on Different Guidance Scale
+Inference Time We compare the inference time of our method with baselines, and the results are compiled in Table 7.
+MEDIC achieves the comparative inference time with generation models and inversion techniques.
+We will make an attempt to reduce the inference time of zero-shot music editing in our future work.
+Method Inference Time
+AudioLDM 2 42.5s
+MusicGen 83.3s
+SDEdit 44.3s
+DDIM Inversion 81.6s
+MusicMagus 89.0s
+DDPM-Friendly 33.3s
+MEDIC 92.0s
+Table 7: Inference Time across different methods.
 Appendix E Potential Negative Societal Impacts
 MEDIC may also lead to potential negative societal impacts that are worthy of consideration.
-If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MEDIC could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation.Hopefully , all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
+If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MEDIC could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation.
+Hopefully, all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
 Appendix F Limitations
-In spite of remarkable outcome of our method , but due to the limitation of the generation model we used , we are incapable of instigating a profound change.
-Though our method is faster than est-optimization and end-to-end diffusion models,but due to the numerous steps it requires (T=200), the duration of computing distance is quite long.Thus , we will implement a more powerful text-to-music generation model to support better editing, while trying to use consistency model or flow matching model to achieve high quality and fast music generation in future work
+In spite of the remarkable outcome of our method, due to the limitation of the generation model we used, we are incapable of instigating a profound change.
+Due to the numerous steps it requires (T=200), the duration of computing distance is quite long.
+Thus, we will implement a more powerful text-to-music generation model to support better editing, while trying to use a consistency model or flow-matching model to achieve high-quality and fast music generation in future work.
+We will make an attempt to edit more interesting and complex music tasks in the future.
 Appendix G Qualitative Results
@@
 G.1 Change Instrument
-In the following figures , we show the capability of MEDIC to change the instrument.Here we edit the ground truth music piece with original prompt "a live recording of ambient acoustic [guitar] music" and editing prompt "a live recording of ambient acoustic [violin] music".
-The difference of instruments can be observed in the Mel-spectrum
+In Figure 5, we show the capability of MEDIC to change the instrument.
+Here we edit the ground truth music piece with the source prompt “a live recording of ambient acoustic [guitar] music" and editing prompt “a live recording of ambient acoustic [violin] music".
+The difference in instruments can be observed in the Mel-spectrum.
 source: a live recording of ambient acoustic [guitar] music
@@
 G.2 Add Instrument
-In the following figures , we show the capability of MEDIC to add more instrument.Here we edit the ground truth music piece with original prompt "a heavy metal instructional audio with a distortion guitar" and editing prompt "a heavy metal instructional audio with a distortion guitar [and drums]".
-The appearance of new instrument can be observed in the Mel-spectrum which present drum sound of high frequency.
+In Figure 6, we show the capability of MEDIC to add more instruments.
+Here we edit the ground truth music piece with the source prompt “a heavy metal instructional audio with a distortion guitar" and editing prompt “a heavy metal instructional audio with a distortion guitar [and drums]".
+The appearance of the new instrument can be observed in the Mel-spectrum which presents a drum sound of high frequency.
 source: a heavy metal instructional audio with a distortion guitar
 edit: a heavy metal instructional audio with a distortion guitar [and drums]
-Figure 6: Editing Type 1 :Add Instrument
+Figure 6: Editing Type 1 Add Instrument
 G.3 Delete Instrument
-In the following figures , we show the capability of MEDIC to delete instrument.Here we edit the ground truth music piece with original prompt "a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and editing prompt "a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood".
-The vanishing of instrument can be observed in the Mel-spectrum
+In Figure 7, we show the capability of MEDIC to delete instruments.
+Here we edit the ground truth music piece with the source prompt “a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and the editing prompt “a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood".
+The vanishing of the instrument can be observed in the Mel-spectrum.
 source: a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood
 edit: a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood
-Figure 7: Editing Type 2 :Delete Instrument
+Figure 7: Editing Type 2 Delete Instrument
 G.4 Change Genre
-In the following figures , we show the capability of MEDIC to change genre of a music piece.
-Here we edit the ground truth music piece with original prompt "a recording of a solo electric guitar playing [blues] licks" and editing prompt "a recording of a solo electric guitar playing [rock] licks".
-The obvious difference of genre can be observed in the Mel-spectrum
+In Figure 8, we show the capability of MEDIC to change the genre of a music piece.
+Here we edit the ground truth music piece with the source prompt “a recording of a solo electric guitar playing [blues] licks" and the editing prompt “a recording of a solo electric guitar playing [rock] licks".
+The obvious difference in genre can be observed in the Mel-spectrum.
 source: a recording of a solo electric guitar playing [blues] licks
 edit: a recording of a solo electric guitar playing [rock] licks
-Figure 8: Editing Type 3 :Change Genre
+Figure 8: Editing Type 3 Change Genre
 G.5 Change Mood
-Mood is an important attribute of a music.
-In the following figures ,we show the capability of MEDIC to change mood of a music piece.
-Here we edit the ground truth music piece with original prompt "a recording of [aggressive] electronic and video game music with synthesizer arrangements"and editing prompt "a recording of [peaceful] electronic and video game music with synthesizer arrangements".
-The change of Mood can be observed in the Mel-spectrum
+Mood is an important attribute of music.
+In Figure 9, we show the capability of MEDIC to change the mood of a music piece.
+Here we edit the ground truth music piece with the source prompt “a recording of [aggressive] electronic and video game music with synthesizer arrangements" and editing prompt “a recording of [peaceful] electronic and video game music with synthesizer arrangements".
+The change of mood can be observed in the Mel-spectrum.
 source: a recording of [aggressive] electronic and video game music with synthesizer arrangements
 edit: a recording of [peaceful] electronic and video game music with synthesizer arrangements
-Figure 9: Editing Type 4 :Change Mood
+Figure 9: Editing Type 4 Change Mood
 G.6 Change Rhythm
 Rhythm represents the speed of the music.
-In the following figures ,we show the capability of MEDIC to change Rhythm of a music piece.
-Here we edit the ground truth music piece with original prompt "a [slow] tempo ukelele tuning recording with static"and editing prompt "a [fast] tempo ukelele tuning recording with static".
-The change of Rhythm can be observed in the Mel-spectrum.The edited Mel-spectrum is much intensive.
+In Figure 10, we show the capability of MEDIC to change the Rhythm of a music piece.
+Here we edit the ground truth music piece with the source prompt “a [slow] tempo ukelele tuning recording with static" and the editing prompt “a [fast] tempo ukelele tuning recording with static".
+The change of Rhythm can be observed in the Mel-spectrum.
+The edited Mel-spectrum is much more intensive.
 source: a [slow] tempo ukelele tuning recording with static
 edit: a [fast] tempo ukelele tuning recording with static
-Figure 10: Editing Type 5 :Change Rhythm
+Figure 10: Editing Type 5 Change Rhythm
 G.7 Change Background
-In the following figures ,we show the capability of MEDIC to change background of instrument of a music piece.
-Here we edit the ground truth music piece with original prompt "an amateur ukulele recording with a [medium to uptempo] pace"and editing prompt "an amateur ukulele recording with a [steady and rhythmic] pace".
+In Figure 11, we show the capability of MEDIC to change the background of the instrument of a music piece.
+Here we edit the ground truth music piece with the source prompt “an amateur ukulele recording with a [medium to uptempo] pace" and editing prompt “an amateur ukulele recording with a [steady and rhythmic] pace".
 The change of instrument background can be observed in the Mel-spectrum.
@@
 edit: an amateur ukulele recording with a [steady and rhythmic] pace
-Figure 11: Editing Type 6 :Change Background
+Figure 11: Editing Type 6 Change Background
 G.8 Change Melody
-In the following figures ,we show the capability of MEDIC to change melody of a music piece.
-Here we edit the ground truth music piece with original prompt "this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo"and editing prompt "this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo".
-The change of Melody can be observed in the Mel-spectrum
+In Figure 12, we show the capability of MEDIC to change the melody of a music piece.
+Here we edit the ground truth music piece with the source prompt “this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo" and editing prompt “this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo".
+The change of Melody can be observed in the Mel-spectrum.
 source: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
 edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
-Figure 12: Editing Type 7 :Change Melody
+Figure 12: Editing Type 7 Change Melody
 G.9 Extract Instrument
-In the following figures ,we show the capability of MEDIC to extract one certain instrument of a music piece.
-Here we edit the ground truth music piece with original prompt "a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]"and editing prompt "a reggae rhythm recording with bongos".
+In Figure 13, we show the capability of MEDIC to extract one certain instrument of a music piece.
+Here we edit the ground truth music piece with the source prompt “a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]" and editing prompt “a reggae rhythm recording with bongos".
 The change of instruments can be observed in the Mel-spectrum.
-Much sounds vanished.
 source: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
 edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
-Figure 13: Editing Type 8 :Extract Instrument
+Figure 13: Editing Type 8 Extract Instrument
 Appendix H safeguards
 In the processing of the data and models involved in this study, we fully considered the potential risks.
-We ensure that all data sources are rigorously screened and vetted,and the model we used is absolutely trained from the safe dataset to minimize security risks of being misused.
+We ensure that all data sources are rigorously screened and vetted, and the model we used is absolutely trained from the safe dataset to minimize the security risks of being misused.

### v4

@@
 MEDIC: Zero-shot Music Editing with Disentangled Inversion Control
-Abstract
-Text-guided diffusion models make a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts.
-Recent works introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pretrained diffusion models for audio modification.
-Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy.
-Moreover, existing editing methods fail to achieve effective complex non-rigid music editing while maintaining essential content preservation and high editing fidelity.
-To counteract these issues, we introduce the Disentangled Inversion technique to disentangle the diffusion process into triple branches, rectifying the deviated path of the source branch caused by DDIM inversion.
-In addition, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention control and cross-attention control with an intermediate Harmonic Branch to progressively achieve the desired harmonic and melodic information in the target music.
-Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing while safeguarding content integrity.
-To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across ten distinct editing categories.
-This facilitates both zero-shot and instruction-based music editing tasks.
-Our method achieves unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques.
-[^1] The project webpage is available at https://MEDIC-Edit.github.io/.
-1 Introduction
-Text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have made great progress in audio generation (Evans et al., 2024a, b), leveraging their impressive capability for realistic and varied outputs.
-In particular, these models (Liu et al., 2023c; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes for specific textual prompts.
-Early audio editing strategies rely on training models from scratch (Copet et al., 2023; Agostinelli et al., 2023) or test-time optimization (Paissan et al., 2023; Plitsis et al., 2024), hampered by intensive computational demands.
-Recent works (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) and Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) inversion techniques, but challenges remain.
-Key among these challenges are maintaining fidelity of editing - ensuring the editing aligns with the provided instructions - and essential content preservation, ensuring that the unaltered musical attributes in the target prompts remain unchanged.
-However, balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving sub-optimal for conditional diffusion models (Mokady et al., 2023).
-Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space.
-However, this comes at the expense of reduced modification capabilities due to noise reduction.
-In this work, we rigorously examine the shortcomings of the DDIM inversion approach.
-Our comprehensive analysis indicates that while techniques like DDIM inversion provide an editable foundation for audio synthesis, they lack precision and may compromise the integrity of the original audio.
-The primary issue stems from the assumption of perfect reversibility in the ordinary differential equation (ODE) process, which is frequently not met during text-conditional editing.
-Consequently, this leads to distortions during the inversion.
-And the implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) aims to improve text adherence.
-However, it inadvertently amplifies the accumulated errors from the inversion process.
-Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has shown promise in achieving high fidelity and essential content preservation.
-For instance, MusicMagus (Zhang et al., 2024) introduces Cross-Attention Control for fine-grained music manipulation of rigid tasks.
-Nevertheless, these methods fail to resolve the issues of accumulated errors and struggle to achieve accurate editing for both rigid and non-rigid tasks, as illustrated in Figure 1.
-We introduce an innovative Disentangled Inversion Control technique to bridge this gap.
-This technique has two main components: Harmonic Attention Control and Disentangled Inversion.
-Cross-attention control (Hertz et al., 2022) and mutual self-attention control (Cao et al., 2023) have demonstrated robust editing capabilities for rigid and non-rigid image editing tasks, respectively.
-However, blindly combining these two approaches sequentially for music editing can result in sub-optimal outcomes, particularly in the original dual-branch setup, where it struggles with global attention refinement.
-To address this challenge, we introduce an intermediate branch called the Harmonic Branch, designed to modify both rigid and non-rigid attributes in music progressively.
-Furthermore, we disentangle the diffusion process into triple branches, correcting the deviation path caused by CFG in the source branch, which affects the essential content preservation.
-The other branches remain unchanged to ensure the highest possible edit fidelity.
-Due to a lack of standardized benchmarks in audio editing, we introduce ZoME-Bench.
-The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories across rigid and non-rigid tasks.
-Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing.
-Experimental results on ZoME-Bench indicate that MEDIC outperforms baselines, achieving significant improvements in essential content preservation and editing fidelity.
-Additionally, MEDIC demonstrates state-of-the-art performance in the variable-length music editing settings of the MusicDelta dataset.
+Abstract.
+Text-guided diffusion models revolutionize audio generation by adapting source audio to specific text prompts.
+However, existing zero-shot audio editing methods such as DDIM inversion accumulate errors across diffusion steps, reducing the effectiveness.
+Moreover, existing editing methods struggle with conducting complex non-rigid music edits while maintaining content integrity and high fidelity.
+To address these challenges, we propose MEDIC, a novel zero-shot music editing system based on innovative Disentangled Inversion Control (DIC) technique, which comprises Harmonized Attention Control and Disentangled Inversion.
+Disentangled Inversion disentangles the diffusion process into triple branches to rectify the deviated path of the source branch caused by DDIM inversion.
+Harmonized Attention Control unifies the mutual self-attention control and the cross-attention control with an intermediate Harmonic Branch to progressively generate the desired harmonic and melodic information in the target music.
+We also introduce ZoME-Bench, a comprehensive music editing benchmark with 1,100 samples covering ten distinct editing categories.
+ZoME-Bench facilitates both zero-shot and instruction-based music editing tasks.
+Our method outperforms state-of-the-art inversion techniques in editing fidelity and content preservation.
+The code and benchmark will be released.
+Audio samples are available at https://melody-edit.github.io/.
+Zero-shot Music Editing, Inversion Techniques, Diffusion Models
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755377isbn: 979-8-4007-2035-2/2025/10ccs: Applied computing Sound and music computingccs: Computing methodologies Natural language generation
+{subcaptionblock}
+0.45 Refer to caption {subcaptionblock}0.45 Refer to caption
+Figure 1. Spider chart for CLAP score comparisons.
+Figure 2. Spider chart for FAD score comparisons.
+Figure 3.
+Comprehensive zero-shot music editing performance evaluation on the ZoME-bench.
+We present spider charts of CLAP scores (higher is better) corresponding to editing fidelity and FAD scores (lower is better) corresponding to source-anchored attribute consistency across 10 editing tasks (e.g., change genre) for DDPM-Friendly (manor2024zero), DDIM Inversion (song2020denoising), MusicMagus (zhang2024musicmagus), and our MEDIC with Disentangled Inversion Control.
+1. Introduction
+Text-guided diffusion models (song2020denoising; song2020improved; peebles2023scalable) have made great progress in audio generation (evans2024fast; evans2024long), leveraging their impressive capability of generating realistic and varied outputs.
+These models (liu2023audioldm2; huang2023makeanaudio; liu2024audiolcm) provide the foundation for prompt-based music editing, offering new opportunities to modify audio landscapes for specific text prompts.
+Early music editing strategies rely on training models from scratch (copet2023simple; agostinelli2023musiclm) or test-time optimization (paissan2023audio; plitsis2024investigating), hence they are hampered by intensive computational demands.
+Recent works (manor2024zero; zhang2024musicmagus) have advanced zero-shot music editing through Denoising Diffusion Implicit Models (DDIM) (song2020denoising) and Denoising Diffusion Probabilistic Models (DDPM) (ho2020denoising) inversion techniques, but challenges remain.
+There are two objectives for music editing, that is, edit fidelity - ensuring the editing aligns with the provided instructions - and essential content preservation - maintaining specified musical properties from the source prompt while modifying only designated attributes.
+For instance, when transforming a melancholic piano composition into an electric guitar arrangement, edit fidelity demands accurate instrument substitution while essential content preservation ensures sustained emotional resonance through retained tempo and recurring minor chord sequences.
+Balancing these two objectives poses a great challenge since it involves a careful exchange of information between the source and target branch in diffusion processes; however, existing inversion methods such as DDIM prove sub-optimal for conditional diffusion models to address this challenge (mokady2023null).
+Enhanced versions of edit-friendly DDPM inversion (hubermanspiegelglas2024edit) make strides in content preservation by imprinting the source onto the noise space.
+However, this method comes at the expense of reduced modification capabilities due to noise reduction.
+In this work, we first examine the shortcomings of the DDIM inversion approach.
+Our comprehensive analysis indicates that while techniques such as DDIM inversion provide a foundation for audio editing, they lack precision and may compromise the integrity of the original audio.
+The primary issue stems from the assumption of perfect reversibility in the ordinary differential equation (ODE) process, which is frequently violated during text-conditional editing.
+This issue leads to distortions during the inversion.
+Although the implementation of Classifier-free Guidance (CFG) (classifier-free) aims to improve text adherence, CFG inadvertently amplifies the accumulated errors from the inversion process.
+Recently, attention control (cao2023masactrl; hertz2022prompt) has shown promise in achieving high fidelity and essential content preservation.
+For instance, MusicMagus (zhang2024musicmagus) introduces Cross-Attention Control for fine-grained music manipulation of rigid tasks.
+Note that in this paper, rigid music editing refers to structural modifications requiring recomposition of core musical relationships, such as the instruments changes, genre transformation, macro-harmonic restructuring, etc; whereas, non-rigid music editing involves parametric adjustments preserving original musical relationships, including adjusting the beat, melody, pitch, rhythm, and other subtle aspects, which generally involves more microscopic editing.
+Nevertheless, attention control methods fail to resolve the issues of accumulated errors and struggle to achieve accurate editing for both rigid and non-rigid tasks, as illustrated in Figure 4.
+In this work, to bridge this gap, we introduce a zero-shot music editing system MEDIC based on an innovative Disentangled Inversion Control technique, which comprises two components of Harmonized Attention Control and Disentangled Inversion.
+Cross-attention control (hertz2022prompt) and mutual self-attention control (cao2023masactrl) have demonstrated robust editing capabilities for rigid and non-rigid image editing tasks, respectively.
+However, simply combining these two approaches sequentially for music editing can result in sub-optimal performance, particularly in the original dual-branch setup, where it struggles with global attention refinement.
+To address this issue, we propose Harmonized Attention Control, which unifies cross-attention control and mutual self-attention control by introducing an intermediate branch named Harmonic Branch, designed to progressively modify both rigid and non-rigid attributes in music.
+Furthermore, we disentangle the diffusion process into triple branches and correct the deviation path caused by CFG in the source branch, which affects the essential content preservation.
+The other branches remain unchanged to ensure high edit fidelity.
+Due to the lack of standardized benchmarks in music editing, we also introduce a new benchmark ZoME-Bench, consisting of 1,100 audio clips in 10 rigorously curated editing categories across rigid and non-rigid tasks.
+Each entry is carefully assembled, comprising a source prompt, a source audio, a target text prompt, human instruction, and blended words intended for editing.
+We then extensively evaluate MEDIC and baselines on ZoME-Bench and other datasets.
 Our contributions can be summarized as follows.
-1) We introduce a novel, training-free methodology called Disentangled Inversion Control (DIC), designed to facilitate consistent manipulations of musical elements and intricate non-rigid editing tasks.
-2) An Harmonized Attention Control framework is introduced to unify cross-attention and mutual self-attention control, which enables both rigid and non-rigid editing.
-3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error by branch disentanglement and correction, aiding in accurately editing the music while preserving the content information.
-4) We build a new benchmark for music editing, named ZoME-Bench, which supports both zero-shot and instruction-based music editing.
-2 Related Works
+  • We introduce a novel, training-free methodology called Disentangled Inversion Control (DIC), designed to facilitate consistent manipulations of musical elements and intricate non-rigid editing tasks.
+We develop a zero-shot music editing system MEDIC based on DIC.
+  • Disentangled Inversion Control includes two critical algorithmic designs.
+(a) A Harmonized Attention Control framework is introduced to unify cross-attention and mutual self-attention control, which enables both rigid and non-rigid editing.
+(b) Disentangled Inversion Technique is proposed to achieve superior results with negligible inversion error by branch disentanglement and correction, aiding in accurately editing the music while preserving the content information.
+  • We introduce ZoME-Bench, a new benchmark for music editing complete with comprehensive evaluation metrics.
+It consists of 1,100 audio clips categorized into 10 rigorously curated editing tasks, encompassing both rigid and non-rigid tasks.
+  • Experimental results on ZoME-Bench indicate that MEDIC outperforms competitive baselines, achieving significant improvements in edit fidelity and essential content preservation, as depicted in Figure 3.
+Moreover, MEDIC achieves state-of-the-art performance under the variable-length music editing settings of the commonly used MusicDelta dataset.
+2. Related Work
 Text-based Audio editing.
-The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt.
-Existing methodologies for addressing these intricate challenges typically follow one of three paths.
-The first involves attempts to develop end-to-end editing models (Copet et al., 2023; Agostinelli et al., 2023; Chen et al., 2024)that employ diffusion processes.
+Some prior text-based audio editing studies utilize diffusion models to manipulate audio content according to the target prompt provided (paissan2023audio; plitsis2024investigating; han2023instructme).
+The two primary challenges (i.e.
+editing fidelity and essential content preservation) command intense focus.
+Existing methodologies (novack2024ditto; wu2024music) for addressing these intricate challenges typically follow one of three paths.
+The first involves attempts to develop end-to-end editing models (copet2023simple; agostinelli2023musiclm; chen2024musicldm) that employ diffusion processes.
 However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets.
-The second path involves test-time optimization strategies that utilize large pre-trained models for editing (Paissan et al., 2023; Plitsis et al., 2024).
-Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text-embeddings for signal reconstruction.
-Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load.
-The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020; Zhang et al., 2024) inversion strategies to extract diffusion noise vectors that match the source signal.
-Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework.
-In this work, we propose a new inversion technique named Disentangled Inversion Control.
-This technique aims to achieve accurate editing while preserving structural information.
+The second path involves test-time optimization strategies that utilize large pre-trained models for editing (paissan2023audio; plitsis2024investigating).
+Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text embeddings for signal reconstruction.
+Some methods choose to employ both strategies (kawar2023imagic), which further increases the computational load.
+The third path involves inversion techniques, which typically use DDPM (hubermanspiegelglas2024edit; wu2023latent) or DDIM (song2020denoising; zhang2024musicmagus) inversion strategies to extract diffusion noise vectors that match the source signal.
+Considering its rapid and intuitive zero-shot editing capabilities, in this work, we choose inversion techniques as our primary research framework.
+Different from existing inversion strategies, we propose a new inversion technique named Disentangled Inversion Control, which disentangles the diffusion process into triple branches with both mutual self-attention control and cross-attention control to achieve accurate editing while preserving structural information.
 Inversion Techniques.
-The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021).
-While the DDIM inversion proves effective for unconditional diffusion models (Song et al., 2020; Wallace et al., 2023), its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing.
-A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges.
+The field of image inversion techniques has experienced significant progress in recent years (brooks2023instructpix2pix; kim2022diffusionclip; parmar2023zero; dhariwal2021diffusion).
+Although DDIM inversion proves to be effective for unconditional diffusion models (song2020denoising; yang2025emovoice; liu2025speech), its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing.
+Various solutions have been proposed to address these challenges (mokady2023null; tumanyan2022plugandplay).
 For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during editing.
-Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024).
-Optimization-based inversion methods using specific latent variables have recently gained popularity (Ju et al., 2023; Kawar et al., 2023).
-These are designed to minimize accumulated errors stemming from the DDIM inversion.
-Techniques such as Null-Text Inversion (Mokady et al., 2023) are promising, but they introduce complexity and instability into the optimization process, making it relatively time-consuming.
-Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
-Figure 1: Comparisons of our method with two-branch inversion techniques, including DDIM Inversion and MusicMagus.
+In contrast, Edit-Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (hubermanspiegelglas2024edit).
+Optimization-based inversion methods using specific latent variables have recently gained popularity (ju2023direct; kawar2023imagic).
+These methods are designed to minimize the accumulated errors that result from the inversion of DDIM.
+Techniques such as Null-Text Inversion (mokady2023null) are promising, but introduce complexity and instability into the optimization process.
+Different from these inversion techniques, we introduce a plug-and-play method called Disentangled Inversion Control to separate branches which enables each branch to unleash its maximum potential individually, achieving superior performance with considerably fewer computational resources.
+Figure 4.
+Comparisons between our method and major baselines based on two-branch inversion techniques, including DDIM Inversion (song2020denoising) and MusicMagus (zhang2024musicmagus).
 (a) Framework of DDIM Inversion, showing configurations with and without classifier-free guidance.
@@
 (c) Framework of our method, featuring Disentangled Inversion Control.
-(d) An illustration comparing the output of the two-branch techniques with the progressive output of our triple-branch approach.
-3 Disentangled Inversion Control
-3.1 Problem Definition and BenchMark Construction
-Despite significant work in text-to-audio generation models, particularly with the emergence of latent diffusion models (LDM), research on zero-shot music editing remains limited.
-Zero-shot music editing seeks to leverage the capabilities of text-to-music generation models to synthesize the desired music, denoted as ${\bm{x}}^{tgt}_{0}$.
-This synthesized music should align with the target edited text prompt ${\mathcal{P}}^{*}$, which is directly modified from the source music ${\bm{x}}^{src}_{0}$ and its corresponding text prompt ${\mathcal{P}}$.
-We compress source audio signals into latent ${\bm{z}}^{src}_{0}$ for inversion.
-To systematically validate our proposed method as a plug-and-play strategy for editing models, compare our method with existing zero-shot music editing methods, and compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset named ZoME-Bench (Zero-shot Music Editing Benchmark).
-ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps (Doh et al., 2023), spanning ten different editing types that include both rigid and non-rigid modifications.
-Each sample is accompanied by its corresponding source prompt, target prompt, human instruction, and source audio.
-Additionally, we include annotations relevant to attention control, such as blended words, to facilitate detailed evaluations. Further details about our benchmark can be found in Appendix B.
-3.2 Motivation
-Figure 1 and Preliminaries in Appendix A reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision.
-This potentially compromising the essential content preservation.
+(d) An illustration comparing the output of the two-branch techniques with the progressive output of our triple-branch method.
+3. Preliminaries And Analyses
+This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied to diffusion models for text-guided audio synthesis.
+We further analyze the challenges associated with these methods.
+3.1. Diffusion Models
+Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition $\bm{c}$} to an output audio ${\bm{z}}_{0}$, corresponding to the given conditioning prompt.
+We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the Gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following this objective:
+[…]
+where noise is added to the sampled data ${\bm{z}}_{0}$ according to diffusion time step $T$.
+During inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using a pre-trained diffusion model for $T$ steps.
+To generate audio from given ${\bm{z}}_{T}$, we employ the deterministic DDIM sampling, where $\alpha$ is hyperparameter:
+(2) ${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$
+3.2. DDIM Inversion
+While diffusion models have superior characteristics in the feature space that can support various downstream tasks, it is hard to apply them to audio in the absence of natural diffusion feature space for non-generated audio.
+Thus, a simple inversion technique known as DDIM inversion is commonly used for unconditional diffusion models, predicated on the presumption that the ODE process can be reversed in the limit of infinitesimally small steps:
+(3) ${\bm{z}}^{*}_{t}=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}{\bm{z}}^{*}_{t-1}+\sqrt{\alpha_{t}}(\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1})\epsilon_{\theta}({\bm{z}}^{*}_{t-1},t-1)$
+where $\bm{z}$$^{*}$ denotes DDIM inversion latent.
+However, in most text-based diffusion models, this presumption cannot be guaranteed, resulting in a perturbation from ${\bm{z}}_{t}$ to ${\bm{z}}^{*}_{t}$ in Equation 2, Equation 3 and Figure 4(a).
+Consequently, an additional perturbation from ${\bm{z}}^{*}_{t}$ to ${\bm{z}}^{src}_{t}$ arises when sampling an audio from ${\bm{z}}^{*}_{T}$ as shown in Figure 4(a):
+(4) ${\bm{z}}_{t}=\sqrt{\alpha_{t}}{\bm{z}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$
+3.3. Classifier-free Guidance
+Classifier-free Guidance (CFG) (classifier-free) is proposed to overcome the limitation of weak text adherence in text-conditioned models. The modified noise estimation in CFG can be expressed as:
+(5) $\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$
+where $\varnothing$ is the embedding of a null text.
+A higher guidance scale $\omega$, which is intended to strengthen the model’s fidelity to the text prompt, inadvertently magnifies the accumulated inversion error.
+CFG further leads to another perturbation from ${\bm{z}}^{src^{\prime}}_{t}$ to ${\bm{z}}^{src^{\prime\prime}}_{t}$ due to the destruction of the DDIM process and error augmentation, as depicted in Figure 4(a).
+This issue becomes problematic in editing scenarios where precise control over audio synthesis is desired.
+4. Methodology
+4.1. Task Definition
+Despite significant work in text-to-audio generation models (huang2023makeanaudio; liu2024audiolcm; liu2023audioldm2; liu2024flashaudio), particularly with the emergence of Latent Diffusion Models (LDM) (liu2023audioldm; chen2024musicldm), research on zero-shot music editing remains limited.
+Given the source music ${\bm{x}}^{src}_{0}$ and its corresponding text prompt $\mathcal{P}$}, zero-shot music editing seeks to leverage the capabilities of text-to-music generation models to directly modify ${\bm{x}}^{src}_{0}$ and $\mathcal{P}$}, and synthesize the desired music ${\bm{x}}^{tgt}_{0}$, which is aligned with the target edited text prompt $\mathcal{P}$$^{*}$.
+We compress source audio signal ${\bm{x}}^{src}_{0}$ into latent ${\bm{z}}^{src}_{0}$ for inversion.
+4.2. Disentangled Inversion Control
+Preliminaries and Figure 4 reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising essential content preservation.
 The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
-In the landscape of prompt-based editing (Dong et al., 2023; Kim et al., 2022; Feng et al., 2023), the ability to grasp the subtleties of linguistics and enable more granular cross-modal interactions stands as a formidable challenge.
-Hertz et al.
-(2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$.
-This leads to the development of various attention control techniques that guide the target denoiser network $\hat{\epsilon}_{\theta}$ in the image domain to better align with target prompts.
-Yet, analogous control mechanisms for non-rigid music editing are noticeably limited.
+In the landscape of prompt-based editing (dong2023prompt; kim2022diffusionclip; feng2023trainingfree), grasping linguistic subtleties and enabling more granular cross-modal interactions remains a formidable challenge.
+hertz2022prompt acknowledges that in image editing, the fusion between text and visual modalities occurs within the parameterized noise prediction network $\epsilon_{\theta}$, which leads to the development of various attention control techniques that guide the target denoiser network $\hat{\epsilon}_{\theta}$ in the image domain to better align with target prompts.
+However, similar control mechanisms for non-rigid music editing are noticeably limited.
 Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach to achieve both rigid and non-rigid music editing.
-DIC strategically disentangles the diffusion process as triple branches, allowing each branch to optimize its functionality.
-At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance.
-We will first introduce Harmonized Attention Control in Section 3.3 and discuss Disentangled Inversion in Section 3.4.
-3.3 Harmonized Attention Control Framework
-The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023b).
-At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer.
-The attention mechanism is formalized as:
-$\text{Attention}(Q,K,V)=MV=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (1)
-where $Q$ denotes the query features derived from the audio features, $K$ and $V$ represent the key and value features, and $M$ is the attention map.
-We explore varying semantic transformations of audio content through harmonized attention control strategies — cross-attention control for rigid changes and mutual self-attention control for non-rigid adjustments.
-Finally, we introduce an intermediate branch to host the desired harmonic and melodic information in the target music.
-The framework of harmonized attention control is depicted in Figure 2.
-Figure 2: The framework of Harmonized Attention Control (HAC).
+DIC strategically disentangles the diffusion process as triple branches (i.e.
+source branch, harmonic branch, and target branch), and allows each branch to optimize its functionality.
+At the same time, the DIC strategy leverages our proposed harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit fidelity.
+Next, we first introduce Harmonized Attention Control in Section 4.3 and discuss Disentangled Inversion in Section 4.4.
+4.3. Harmonized Attention Control Framework
+The denoising architecture $\epsilon_{\theta}$ is structured as a sequence of fundamental blocks, each comprising a residual block (he2015deep) followed by self-attention and cross-attention modules (vaswani2023attention; dosovitskiy2020image; liu2023vit).
+The proposed harmonized attention control (HAC) framework is depicted in Figure 5.
+We explore varying semantic transformations of audio content through harmonized attention control strategies - cross-attention control for rigid editings and mutual self-attention control for non-rigid editings.
+We introduce an intermediate Harmonic Branch to host the desired harmonic and melodic information in the target music.
+Below we elaborate on Cross-Attention Control, Mutual Self-Attention Control, and Harmonic Branch Integration within HAC.
+Figure 5.
+The Harmonized Attention Control (HAC) framework.
 HAC unifies cross-attention control and mutual self-attention control with an additional branch named Harmonic Branch to host the desired composition and structural information in the target music.
-3.3.1 Cross-Attention Control
+4.3.1. Cross-Attention Control
 Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt $\mathcal{P}$}, into a second generation with the target prompt $\mathcal{P}$$^{*}$.
-We follow the practice of Prompt-to-Prompt (Hertz et al., 2022) and define CAC as Global Attention Refinement and Local Attention Blend.
+Motivated by Prompt-to-Prompt (hertz2022prompt), We define CAC as Global Attention Refinement and Local Attention Blend.
 Global Attention Refinement
-At a given time step $t$, the attention map $M_{t}$ for both the origin and target branches is computed, averaging over all layers with respect to the noised latent $\mathbf{z}_{t}$.
-We employ an alignment function $A$ that maps each token index from the target prompt $\mathcal{P}^{*}$ to its equivalent in $\mathcal{P}$ or to None for non-aligning tokens.
-The refinement action is thus:
-$\text{Refine}(M^{src}_{t},M^{tgt}_{t},t)=\begin{cases}(M^{tgt}_{t})_{i,j}&\text{if }A(j)=\text{None},\\ (M_{t})_{i,j}&\text{otherwise}.\end{cases}$ (2)
+At a given time step $t$, the attention map $M_{t}$ for both source and target branches is computed, averaging over all layers with respect to the noised latent $\mathbf{z}_{t}$.
+We employ an alignment function $A$ that maps each token index from the target prompt $\mathcal{P}^{*}$ to its equivalent in $\mathcal{P}$, or to None for unaligned tokens.
+The refinement action is:
+(6) $\text{Refine}(M^{src}_{t},M^{tgt}_{t},t)=\begin{cases}(M^{tgt}_{t})_{i,j}&\text{if }A(j)=\text{None},\\ (M_{t})_{i,A(j)}&\text{otherwise}.\end{cases}$
 Local Attention Blends
-Beyond global attention, we incorporate a blending mechanism as suggested by  Hertz et al.
-(2022) and  Mokady et al.
-(2023).
-This method selectively integrates and maintains certain semantics by using target blend words $w^{tgt}$ for semantic additions and source blend words $w^{src}$ for semantic preservation.
+Beyond global attention, we incorporate a blending mechanism suggested by hertz2022prompt and  mokady2023null.
+This method selectively integrates and maintains certain semantics by using target blend words $w^{tgt}$, which are words in the target prompt whose semantics need to be added; and source blend words $w^{src}$, which are words in the source prompt whose semantics need to be preserved.
 At each denoising step $t$, the mechanism operates on the target latent ${\bm{z}}^{tgt}_{t}$ as follows:
-$m_{\text{tgt}}$ $=\text{Threshold}\left[M^{tgt}_{t},w_{\text{tgt}},k_{\text{tgt}}\right],$ (3)
-$m_{\text{src}}$ $=\text{Threshold}\left[M^{src}_{t},w_{\text{src}},k_{\text{src}}\right],$ (4)
-${\bm{z}}^{tgt}_{t}$ $=(1-m^{tgt}+m^{src})\odot{\bm{z}}^{src}_{t}+(m^{tgt}-m^{src})\odot{\bm{z}}^{tgt}_{t}$ (5)
-where $m_{\text{tgt}}$ and $m_{\text{src}}$ are binary masks and threshold function is as delineated below:
-$\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$ (6)
+(7) $m_{\text{tgt}}$ $=\text{Threshold}\left[M^{tgt}_{t}(w_{\text{tgt}}),k_{\text{tgt}}\right],$
+(8) $m_{\text{src}}$ $=\text{Threshold}\left[M^{src}_{t}(w_{\text{src}}),k_{\text{src}}\right],$
+(9) ${\bm{z}}^{tgt}_{t}$ $=(1-m^{tgt}+m^{src})\odot{\bm{z}}^{src}_{t}+(m^{tgt}-m^{src})\odot{\bm{z}}^{tgt}_{t}$
+where $m_{\text{src}}$ and $m_{\text{tgt}}$ are binary masks obtained by calibrating the aggregated attention maps $M^{src}_{t}(w^{src})$ and $M^{tgt}_{t}(w^{tgt})$ with threshold parameters $k_{src}$ and $k_{tgt}$, using the following threshold function:
+(10) $\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$
 For simplicity, we define the process of local editing in Equation 9 as:
-${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$ (7)
+(11) ${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$
 Scheduling Cross-Attention Control
-Applying cross-attention control only at early stages ensures creative flexibility while maintaining structure.
-Acting on insights from  Hertz et al.
-(2022), we limit cross-attention to the initial phases up to a cutoff point $\tau_{c}$.
-This moderation allows us to capture the nuances and intended changes in musical compositions effectively.
-The approach is defined as:
-$\text{CrossEdit}(M^{src},M^{tgt},t)=\begin{cases}\text{Refine}(M^{src}_{t},M^{tgt}_{t})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}_{t}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
-3.3.2 Mutual Self-Attention Control
-We diverge from the conventional use of cross-attention mechanisms and instead draw inspiration from the MasaCtrl (Cao et al., 2023) technique to refine music structure through self-attention queries.
-These queries adeptly navigate through non-rigid musical transformations, aligning with the designated musical theme or instrument (target prompt).
-The process begins by sketching the foundational musical theme using the target’s self-attention components—$Q^{tgt}$, $K^{tgt}$, and $V^{tgt}$.
-This is followed by enriching this theme with elements resembling the thematic content from the source ($K^{src}$, $V^{src}$), steered by $Q^{tgt}$.
-However, applying this attentive modulation uniformly over all processing layers and through every denoising step might result in a composition excessively mirroring the source.
-Consequently, echoing the ethos of MasaCtrl, our proposed solution selectively employs mutual self-attention in the decoder portion of our music editing U-Net, initiated after a set number of denoising iterations.
+Implementing cross-attention control throughout the entire sampling schedule can cause excessive focus on music structures, hindering the ability to incorporate intended changes.
+In contrast, applying it only during the early stages allows for creative flexibility while still preserving structural integrity.
+Therefore, we limit cross-attention to the initial phases up to a cutoff point $\tau_{c}$.
+This moderation allows us to effectively capture the nuances and intended changes in musical compositions.
+The scheduling approach is defined as follows:
+(12) $\text{CrossEdit}(M^{src},M^{tgt},t)=$
+[…]
+4.3.2. Mutual Self-Attention Control
+[…]
 Scheduling Mutual Self-Attention Control
-The application of mutual self-attention is meticulously planned, beginning at a specific denoising step $S$ and extending beyond a designated layer $L$.
-The strength and influence of this control mechanism are designed as follows:
-$\text{SelfEdit}({Q^{src},K^{src},V^{src}},{Q^{tgt},K^{tgt},V^{tgt}},t)=\begin{cases}{Q^{src},K^{src},V^{src}}&\text{if }t\geq S\text{ and }l\geq L,\\ {Q^{tgt},K^{src},V^{src}}&\text{otherwise}\\ \end{cases}$ (9)
-In this framework, $S$ signifies the denoising step from which the mutual self-attention control commences, serving as a temporal threshold.
-Similarly, $L$ distinguishes the layer index below which this nuanced control strategy becomes operational, tailoring the musical output towards the intended artistic direction.
-3.3.3 Harmonic Branch Integretion
-The naive combination of cross-attention control and mutual self-attention control sequentially would lead to sub-optimal results in the original dual-branch setup, especially failing the global attention refinement.
-To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music.
-Our unified framework is detailed in Algorithm 2.
+The application of the nuanced mutual self-attention control is meticulously planned, beginning at a specific denoising step $S$ and extending beyond a designated layer index $L$.
+The strength and influence of this control mechanism are designed as follows.
+(13) $\text{SelfEdit}({Q^{src},K^{src},V^{src}},{Q^{tgt},K^{tgt},V^{tgt}},t)=$
+[…]
+In this framework, $S$ serves as a temporal threshold while $L$ tailors the musical output towards the intended artistic direction.
+4.3.3. Harmonic Branch Integretion
+We hypothesize that a simple sequential combination of cross-attention control and mutual self-attention control may yield sub-optimal results within the original dual-branch framework.
+This approach is particularly ineffective in refining global attention, as depicted in Figure 4(d).
+Our empirical validation, presented in Table 2 (HAC vs w/o HB), supports this hypothesis.
+To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate branch to host the desired composition and structural information in the target music.
+The unified framework of Harmonized Attention Control is detailed in Algorithm 1.
 During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$.
-This lays the formal structure of the target music.
-Following this, cross-attention control is applied on $M^{har}$ and $M^{tgt}$ to refine the semantic information for $M^{tgt}$.
-As illustrated in Figure 2, the harmonic branch output $z^{har}_{0}$ reflects the requested non-rigid changes (e.g., “violin"), while preserving the rigid content semantics (e.g., “with noise").
-The target branch output $z^{tgt}_{0}$ builds upon the structural layout of the $z^{har}$ while reflecting the requested rigid changes (e.g., “with noise").
+This latent lays the formal structure of the target music.
+Next, cross-attention control is applied on $M^{har}$ and $M^{tgt}$ to refine the semantic information for $M^{tgt}$.
+As illustrated in Figure 5, the harmonic branch output $z^{har}_{0}$ reflects the requested non-rigid changes (e.g., “rapid”), while preserving the rigid content semantics (e.g., “guitar”).
+The target branch output $z^{tgt}_{0}$ builds upon the structural layout of the $z^{har}$ while reflecting the requested rigid changes (e.g., “violin”).
 Algorithm 1 Harmonized Attention Control in one DDIM Forward Process
 1:Input: A source prompt $\mathcal{P}$}, a target prompt $\mathcal{P}$$^{*}$, a source audio latent ${\bm{z}}_{0}$, denoising network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$, current time step $\tau$, source and target blend words $w_{\text{src}},w_{\text{tgt}}$, input latents $z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}$.
-2:  $\epsilon_{\text{src}},\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},M_{\text{src}}=\epsilon_{\theta}(z^{\text{src}}_{\tau},\tau,c_{\text{src}})$
-3:  $\epsilon_{\text{tgt}},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},M_{\text{tgt}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}})$
-4:  $\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\}=\text{SelfEdit}(\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},\tau)$
-5:  $\epsilon_{\text{har}},M_{\text{har}}=\epsilon_{\theta}(z^{\text{har}}_{\tau},\tau,c_{\text{src}};\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\})$
-6:  $\hat{M^{\text{tgt}}}=\text{CrossEdit}(M_{\text{har}},M_{\text{tgt}},\tau)$
-7:  $\hat{\epsilon_{\text{tgt}}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}};\hat{M^{\text{tgt}}})$
-8:  $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}=\text{Sample}([z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}],[\epsilon_{\text{src}},\epsilon_{c_{\text{tgt}}},\epsilon_{\text{har}}],\tau)$
-9:  ${\bm{z}}^{tgt}_{\tau-1}$ = $\text{LocalEdit}({\bm{z}}^{src}_{\tau-1},{\bm{z}}^{tgt}_{\tau-1},M^{src}_{\tau-1},M^{tgt}_{t},w_{src},w_{tgt})$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 10:Output: $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}$
-3.4 Disentangled Inversion Technique
-Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal.
-Increasing the guidance scale enhances editability, but sacrifices reconstruction accuracy due to latent code deviation during editing.
-In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the harmonic, and the target branch.
+4.4. Disentangled Inversion Technique
+Using DDIM inversion without classifier-free guidance yields an easily modifiable but imprecise approximation of the original audio signal.
+Increasing the classifier-free guidance scale enhances editability, but sacrifices reconstruction accuracy due to latent code deviation during editing.
+In order to address these limitations, we propose Disentangled Inversion Technique to disentangle the diffusion process into three branches: the source branch, the harmonic branch, and the target branch, with the detailed algorithm outlined in Algorithm 2.
 This decoupling is designed to unleash the capabilities of each branch separately.
@@
 This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio.
-On the other hand, the target branch and harmonic branch are left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio.
-The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio.
-We will further discuss this in the Section 4.2.2.
-The algorithm of Disentangled Inversion Technique has been outlined in Algorithm 2.
-Typical diffusion-based editing (Han et al., 2023b; Miyake et al., 2023) involves two parts: an inversion process to get the diffusion space of the audio, and a forward process to perform editing on the diffusion space.
-Disentangled Inversion can be plug-and-played into the forward process and rectifies the deviation path step by step.
+On the other hand, the target branch and harmonic branch are left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio, thereby ensuring the fidelity and integrity of the generated audio.
+Effectiveness of Disentangled Inversion Technique are verified and discussed in Section 2.
+Typical diffusion-based editing (han2023improving; miyake2023negativeprompt) involves two parts: an inversion process to obtain the diffusion space of the audio, and a forward process to perform editing on the diffusion space.
+In contrast, our Disentangled Inversion Technique can be plug-and-played into the forward process and rectifies the deviation path step by step.
 Specifically, Disentangled Inversion first computes the difference between ${\bm{z}}^{*}_{t-1}$ and ${\bm{z}}^{src}_{t-1}$, then adds the difference back to ${\bm{z}}^{src}_{t-1}$ in DDIM forward.
-We only add the difference of the source prompt in latent space and update ${\bm{z}}^{src}_{t-1}$, which is the key to retaining the editability of the target prompt’s latent space.
+We only add the difference of the source prompt in latent space and update ${\bm{z}}^{src}_{t-1}$, which is the key to retaining the editability of the latent space of the target prompt.
 Algorithm 2 Disentangled Inversion Technique
@@
 4: Initialize ${\bm{z}}^{src}_{T}\leftarrow{\bm{z}}^{*}_{T}$, ${\bm{z}}^{tgt}_{T}\leftarrow{\bm{z}}^{*}_{T}$, ${\bm{z}}^{har}_{T}\leftarrow{\bm{z}}^{*}_{T}$.
-5:  for $t=T$ to $1$ do
-6:     $[{\bm{d}}^{src}_{t-1},{\bm{d}}^{tgt}_{t-1},{\bm{d}}^{har}_{t-1}]\leftarrow{\bm{z}}^{*}_{t-1}$ DDIM_Forward$({\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega)$
-7:     ${\bm{z}}^{src}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
-8:     ${\bm{z}}^{har}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{har}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
-9:     ${\bm{z}}^{tgt}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{tgt}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
+5:for $t=T$ to 11 do
+6:  ${\bm{d}}^{src}_{t-1},{\bm{d}}^{tgt}_{t-1},{\bm{d}}^{har}_{t-1}]\leftarrow{\bm{z}}^{*}_{t-1}$
+[…]
+7:  ${\bm{z}}^{src}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + ${\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
+8:  ${\bm{z}}^{har}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{har}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + ${\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
+9:  ${\bm{z}}^{tgt}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{tgt}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + ${\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
 10:end for
-11:  return  ${\bm{z}}^{tgt}_{0}$
-4 Experiments
-Implementation Details.
-We infer different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023c) models with 200 inference steps.
-The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA A800 GPU for a fair comparison.
-We evaluate all methods in ZoMo-Bench across all editing types with fixed length.
-And we use the MusicDelta subset of the MelodyDB (Bittner et al., 2014) dataset for variable length comparisons, comprised of 34 musical excerpts in varying styles and in lengths ranging from 20 seconds to 5 minutes.
-Details can be found in Appendix C.
-Evaluation Metrics.
-Our models employ a comprehensive evaluation using both objective and subjective metrics to assess essential content preservation, text-audio alignment fidelity, and audio quality.
-Objective metrics include Structure Distance (SD) (Ju et al., 2023), CLAP (Contrastive Language-Audio Pretraining) Score (Elizalde et al., 2023), LPAPS (Learned Perceptual Audio Patch Similarity) (Iashin & Rahtu, 2021; Paissan et al., 2023), and FAD (Fréchet Audio Distance) (Kilgour et al., 2018).
-The CLAP Score evaluates how well the edited audio aligns with the target prompt, while Structure Distance, LPAPS, and FAD, which have been adapted from image domain metrics, measure the similarity between the edited and source audio.
-Specifically, LPAPS quantifies the consistency of the edited audio in relation to the source audio, and the FAD metric assesses the distance between two distributions of audio signals.
-The Structure Distance measures the structural similarity between the edited and source audio.
-For the subjective evaluation, we conduct crowd-sourced human assessments using the Mean Opinion Score (MOS) to evaluate both editing fidelity (MOS-Q) and content preservation (MOS-P).
-We attach the details of all metrics in Appendix C.1.
-4.1 Zero-shot Music Editing Results
+11:Return: ${\bm{z}}^{tgt}_{0}$
+5. Experiments
+Table 1.
+Performance comparison of objective and subjective zero-shot music editing on ZoME-Bench (fixed length) and MusicDelta (variable length).
+The evaluation focuses on content preservation (SD, LPAPS, FAD, Chroma, MOS-P) and edit fidelity (CLAP Score, Accuracy, MOS-Q).
+Results include mean and standard deviation for both edit fidelity (MOS-Q) and content preservation (MOS-P).
+Metrics achieving the best scores are highlighted in bold.
 Method Objective Metrics Subjective Metrics
 […]
-AudioLDM 2 23.86 0.21 10.36 0.58 73.48 70.12
-MusicGen 23.39 0.21 6.63 0.59 75.46 71.28
-SDEdit 25.87 0.22 12.18 0.40 69.38 66.23
-DDIM Inversion 22.52 0.21 9.51 0.49 73.10 3.38
-MusicMagus 16.23 0.19 5.15 0.55 75.12 74.34
-DDPM-Friendly 18.30 0.19 5.16 0.54 75.27 73.86
-MEDIC 11.97 0.15 2.49 0.61 79.81 77.29
-Table 1: Comparison with baselines on ZoME-Bench with fixed length about structure, content preservation, and CLAP similarity.
-We present a comparative study of our Disentangled Inversion Control (DIC) against several established music generation and editing baselines, including AudioLDM 2 (Liu et al., 2023c), MusicGen (Copet et al., 2023), SDEdit (Liu et al., 2023a), DDIM Inversion Ho et al.
-(2020), MusicMagus (Zhang et al., 2024), and DDPM-Friendly (Manor & Michaeli, 2024).
-Utilizing the ZoMo-Bench test set we developed, we assess the quality of the generated audio samples, focusing on two key aspects: Content Preservation and Editing Fidelity.
-The results, summarized in Table 1, lead to the following conclusions: (1) Our method, MEDIC, significantly outperforms both generation-based and inversion-based models in terms of editing fidelity and audio similarity, demonstrating its effectiveness in addressing complex editing tasks.
-(2) Although DDPM-Friendly and MusicMagus show improvements in essential content preservation, they struggle to maintain high text-audio alignment compared to generation models.
-MEDIC effectively addresses accumulated errors, achieving high audio similarity and an impressive CLAP score, reflecting its superior editing fidelity.
-(3) MEDIC also excels in subjective metrics, achieving the best performance in audio quality and content preservation.
-Variable Length Comparisons We further evaluate MEDIC against baseline methods in a variable length setting, with the results presented in Table 2.
-Our analysis leads to the following conclusions: (1) MEDIC outperforms all baselines across all metrics, demonstrating the effectiveness of our approach in variable-length scenarios.
-(2) Inversion-based baselines experience a significant degradation in CLAP score, while MEDIC maintains the highest CLAP score of 0.51 on the MusicDelta dataset.
-Method Objective Metrics Subjective Metrics
+Fixed Length Comparisons on ZoME-Bench
 […]
-AudioLDM 2 24.40 0.22 7.07 0.44 66.37 64.28
-MusicGen 27.71 0.23 7.70 0.46 67.41 63.76
-SDEdit 28.12 0.24 13.21 0.24 62.80 62.18
-DDIM Inversion 23.5 0.21 10.12 0.27 65.94 65.73
-MusicMagus 25.6 0.22 7.13 0.43 67.45 67.12
-DDPM-Friendly 21.53 0.23 6.68 0.30 66.34 67.28
-MEDIC 19.5 0.20 6.58 0.51 71.62 70.18
-Table 2: Comparison with baselines in variable length setting.
-(a) Spider chart for CLAP score comparisons .
-(b) Spider chart for FAD score comparisons.
-Figure 3: A comprehensive performance evaluation on the ZoME-bench.
-We present spider charts of CLAP scores and FAD scores across 10 editing tasks for DDPM-Friendly, DDIM Inversion, MusicMagus, and Disentangled Inversion Control methods.
-Fine-grained Comparisons on ZoME-Bench To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing.
-The FAD and CLAP scores are depicted in Figure 3.
-The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks.
-(2) While baselines exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing non-rigid manipulations, as reflected in their inferior performance, especially in the “Change Genre" and “Change Melody" non-rigid editing types.
-4.2 Ablation Study
-The ablation studies presented in this section aim to validate the contributions of the Harmonized Attention Control and Disentangled Inversion to our framework’s overall performance.
-Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in Appendix D.
-4.2.1 Ablation on Attention Control Methods
 […]
-HAC 11.97 0.15 2.49 0.61
-w/o MSA Control 14.13 0.19 2.75 0.56
-w/o CA Control 13.78 0.18 2.50 0.58
-w/o Harmonic Branch 12.75 0.16 2.59 0.59
-Table 3: Ablation results about different control methods. MSA: Mutual Self-Attention, CA: Cross Attention.
-To validate the impact of our attention control mechanisms, we perform ablations on the following configurations: Remove Mutual Self Attention Control (w/o MSA Control), Remove Cross Attention Control (w/o CA Control), and without Harmonic Branch (w/o Harmonic Branch).
-The findings, detailed in Table 3, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance.
-(2) Although the naive combination of mutual self-attention control and cross-attention control improves the preservation and fidelity, it still yields sub-optimal outcomes due to the lack of a progressive harmonic branch.
-This demonstrates the effectiveness of harmonized attention control.
-4.2.2 Ablation on Disentangled Inversion Technique
-To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from Table 4: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score.
-2) Incorporating target distance into the target branch and harmonic branch improves structural integrity and content preservation.
-However, this modification leads to a significant decrease in the CLAP score, indicating that while structural quality may improve, the alignment with the target prompt suffers.
-3) Adding harmonic distance to the harmonic branch leads to a noticeable decline in performance.
-This finding showcases that introducing excessive deviations in the harmonic branch may adversely affect the overall audio quality and coherence.
 […]
-$[{\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ 14.28 0.17 2.64 4.67 0.57
-$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{src}]$ 13.17 0.17 2.55 4.59 0.58
-$[{\bm{d}}_{src},{\bm{d}}_{tgt},{\bm{0}}]$ 11.24 0.15 2.44 4.32 0.56
-$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 4.29 0.56
-$[{\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$ 37.51 0.27 27.2 14.23 0.28
-$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.61
-Table 4: Ablation results from the disentangled inversion technique.
-[,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance (line 6 in Algorithm 2).
-MSE means the mean square error loss between the edited audio features and source audio features.
-4.3 Qualitative Results
-Figure 4: Visualizations of the source audio’s mel and edited mel-spectrograms by different editing methods.
+[…]
+[…]
+[…]
+[…]
+Variable Length Comparisons On MusicDelta
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+5.1. Experimental Setup
+Datasets.
+To address the absence of a standardized benchmark for inversion and editing techniques and to systematically validate our proposed method as a plug-and-play strategy for music editing and compare with existing zero-shot music editing methods, we construct a new benchmark ZoME-Bench (Zero-shot Music Editing Benchmark).
+ZoME-Bench comprises 1,100 audio samples, which are selected from MusicCaps (agostinelli2023musiclm) ZoME-Bench covers 10 different editing types that include both rigid and non-rigid modifications, detailed in supplementary material.
+Each sample is accompanied by its corresponding source prompt, target prompt, human instruction, and source audio.
+We evaluate different editing methods using the pre-trained AudioLDM 2 model (liu2023audioldm2) with 200 inference steps, on NVIDIA A800 GPU.
+The setting of all baselines follows that of  manor2024zero.
+All baseline editing methods are first evaluated on ZoME-Bench across all editing types of 10 seconds duration.
+We further use the commonly used MusicDelta subset of the MedleyDB (bittner2014medleydb) dataset for variable length performance comparisons.
+MusicDelta comprises 34 musical excerpts in varying styles and in lengths ranging from 20 seconds to 5 minutes.
+Evaluation Metrics.
+Our comprehensive evaluation involves both objective and subjective metrics to assess essential content preservation, text-audio alignment fidelity, and audio quality.
+(1) Objective Metrics: We use a variety of objective metrics to measure different aspects of audio editing effectiveness.
+Structure Distance (SD) (ju2023direct), CLAP Score (Contrastive Language-Audio Pretraining) (elizalde2023clap), LPAPS (Learned Perceptual Audio Patch Similarity) (iashin2021taming; paissan2023audio), Chroma (Chromagram Similarity), FAD (Fréchet Audio Distance) (kilgour2018fad), and Accuracy are included in our evaluation framework.
+(2) Subjective Metrics: For subjective evaluation, we conduct crowd-sourced human assessments using the Mean Opinion Score (MOS).
+This metric is used to evaluate both edit fidelity (MOS-Q) and content preservation (MOS-P).
+5.2. Zero-shot Music Editing Results
+We present a comparative study of our MEDIC with DIC against several established music generation and editing baselines.
+We group these baselines into generation-based methods, including AudioLDM 2 (liu2023audioldm2) and MusicGen (copet2023simple), and inversion-based methods, including SDEdit (liu2023audioldm), DDIM Inversion (ho2020denoising), MusicMagus (zhang2024musicmagus), and DDPM-Friendly (manor2024zero).
+Fixed Length Comparisons.
+We evaluate the generated audio samples on the ZoME-Bench test set with a fixed length, focusing on the key aspects of content preservation and edit fidelity.
+As exhibited in Table 1, the results yield the following insights: (1) Our MEDIC substantially outperforms both generation inversion-based models in terms of edit fidelity and content preservation in both objective and subjective metrics, demonstrating its effectiveness in addressing complex editing tasks.
+(2) While DDPM-Friendly and MusicMagus improve content preservation (higher MOS-P), they lag in text-audio alignment and editing precision, indicated by their lower CLAP Scores and Accuracy.
+In contrast, MEDIC achieves consistently better CLAP, Accuracy, LPAPS, FAD, and Chroma, demonstrating superior alignment with target prompts as well as overall musical similarity.
+Variable Length Comparisons We further evaluate MEDIC against the baseline methods in a variable length setting on the MusicDelta dataset.
+Table 1 highlights the following insights: (1) MEDIC consistently outperforms all baselines across all objective and subjective metrics, demonstrating strong robustness and adaptability for editing longer audio segments.
+(2) Inversion-based baselines suffer a notable drop in edit fidelity, as seen in their lower Accuracy and CLAP scores, likely due to error accumulation and insufficient attention control.
+In contrast, MEDIC achieves the highest CLAP and Accuracy, confirming its ability to deliver precise and well-aligned edits even on audio longer than 20 seconds.
+Fine-grained Comparisons on ZoME-Bench.
+We further compare performance across different editing types using FAD and CLAP scores (Figure 3): (1) MEDIC consistently outperforms all baselines for both rigid and non-rigid editing tasks.
+(2) Baselines can handle some rigid edits, but perform poorly on non-rigid manipulations such as “Change Genre” and “Change Melody”.
+Figure 6. Visualizations of the mel-spectrograms of the source audio and the edited audio by different editing methods.
+5.3. Ablation Study and Visualization
+Effect of Attention Control Methods.
+To validate our attention control designs, we conduct ablation studies on the following configurations: Remove Mutual Self-Attention Control (w/o MSA Control), Remove Cross-Attention Control (w/o CA Control), and Remove Harmonic Branch (w/o Harmonic Branch).
+Table 2 reveals that: (1) Both cross-attention and mutual self-attention controls significantly enhance editing performance.
+Their presence leads to improvements across all metrics, demonstrating their crucial role in achieving higher content preservation and edit fidelity.
+(2) While the naive combination of mutual self-attention and cross-attention control improves metrics such as LPAPS, FAD, and Chroma, it results in sub-optimal outcomes without the inclusion of the Harmonic Branch.
+This missing component reduces overall coherence and refinement, emphasizing the critical role of the Harmonic Branch in augmenting attention control mechanisms to optimize both content integrity and edit fidelity.
+Table 2. Effect of Mutual Self-Attention control (MSA), Cross-Attention (CA) control, and Harmonic Branch (HB).
+[…]
+HAC 11.97 0.15 2.49 0.73 0.61 0.59
+w/o MSA Ctrl 14.13 0.19 2.75 0.57 0.56 0.45
+w/o CA Ctrl 13.78 0.18 2.50 0.63 0.58 0.48
+w/o HB 12.75 0.16 2.59 0.66 0.59 0.51
+Effectiveness of Disentangled Inversion Technique.
+We assess the soundness of Algorithm 2 and the effectiveness of disentanglement of triple branches.
+As shown in Table 3, 1) When employing ${\bm{d}}_{src},{\bm{0}},{\bm{0}}]$—the configuration used in MEDIC—results are notably strong across several metrics.
+MEDIC achieves the highest Accuracy (0.59), CLAP score (0.61), and Chroma (0.73), demonstrating its balanced approach in effectively preserving content fidelity while maintaining strong alignment with the target prompt.
+2) Configurations that incorporate additional distances, such as ${\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ and ${\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$, show a notable decline in performance across most metrics.
+3)Interestingly, the configuration ${\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ achieves the lowest values for LPAPS and FAD, suggesting potential improvements in audio similarity.
+However, this setup seems to markedly reduce overall text-audio alignment and accuracy, indicating a trade-off where structural precision comes at the expense of editorial coherence.
+Table 3.
+Ablation study of Disentangled Inversion Technique.
+$\cdot$$\cdot,\cdot$] denotes adding the distance (line 6 in Algorithm 2).
+MSE is the mean square error loss between the edited audio features and source audio features.
+${\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ is used in MEDIC.
+[…]
+${\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 0.73 0.61 0.59
+${\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ 14.28 0.17 2.64 0.53 0.57 0.47
+${\bm{d}}_{src},{\bm{0}},{\bm{d}}_{src}]$ 13.17 0.17 2.55 0.47 0.58 0.52
+${\bm{d}}_{src},{\bm{d}}_{tgt},{\bm{0}}]$ 11.24 0.15 2.44 0.52 0.56 0.52
+${\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 0.55 0.56 0.32
+${\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$ 37.51 0.27 2.72 0.32 0.28 0.31
+${\bm{0}},{\bm{0}},{\bm{0}}]$ 12.65 0.16 2.54 0.47 0.57 0.47
+Visualization.
 To complement our quantitative findings, we present a qualitative comparison in Figure 6.
-Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with preserving melodic content and harmonic structure.
-In contrast, our Disentangled Inversion Control performs better in precise music editing while preserving structural integrity.
-We provide additional qualitative results in the Appendix G for further examination.
-5 Conclusion
-In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts.
-We proposed the Disentangled Inversion Control to support both rigid and non-rigid editing tasks.
-Instead of a two-branch setting, we add an intermediate branch named the harmonic branch to progressively integrate harmonic and melodic information in music by cross-attention control and mutual self-attention control.
-To counteract the accumulated errors caused by DDIM inversion and CFG, we introduced a simple but effective method named disentangled inversion to separate the diffusion process into triple branches and eliminate the latent discrepancy distance in the source branch.
-Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,100 samples across 10 varied editing categories—attested to the superiority of our methods.
-And the experiment results on variable length test sets and ablation studies further validated the effectiveness and robustness of our method.
+Methods such as SDEdit and inversion-based techniques often struggle to balance high editability and preserve melodic content and harmonic structure.
+In contrast, MEDIC performs better in precise music editing while preserving structural integrity.
+We provide additional qualitative results for all editing categories in supplementary material, demonstrating the superiority of our approach.
+6. Conclusion
+In this paper, we propose the Disentangled Inversion Control to support both rigid and non-rigid editing tasks and develop a zero-shot music editing framework MelodyEdit based on DIC.
+We add an intermediate harmonic branch to progressively integrate harmonic and melodic information in music by cross-attention control and mutual self-attention control.
+To counteract the accumulated errors caused by DDIM inversion and CFG, we introduce Disentangled Inversion to separate the diffusion process into triple branches and eliminate the latent discrepancy distance in the source branch.
+Extensive experiments demonstrate superiority of MelodyEdit on both fixed and variable length settings.
 We envisage that our work could serve as a basis for future zero-shot music editing studies.
-Appendix A Preliminaries And Analyses
-This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied to diffusion models for text-guided audio synthesis.
-It further explores the challenges associated with these methods.
-A.1 Diffusion Models
-Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, corresponding to the given conditioning prompt.
-We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the Gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
-[…]
-Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$.
-At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using a pre-trained network for $T$ steps.
-To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM sampling:
-${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (11)
-A.2 DDIM Inversion
-While diffusion models have superior characteristics in the feature space that can support various down-stream tasks, it is hard to apply them to audios in the absence of natural diffusion feature space for non-generated audios.
-Thus, a simple inversion technique known as DDIM inversion is commonly used for unconditional diffusion models, predicated on the presumption that the ODE process can be reversed in the limit of infinitesimally small steps:
-${\bm{z}}^{*}_{t}=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}{\bm{z}}^{*}_{t-1}+\sqrt{\alpha_{t}}(\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1})\epsilon_{\theta}({\bm{z}}^{*}_{t-1},t-1)$ (12)
-However, in most text-based diffusion models, this presumption cannot be guaranteed, resulting in a perturbation from ${\bm{z}}_{t}$ to ${\bm{z}}^{*}_{t}$.
-Consequently, an additional perturbation from ${\bm{z}}^{*}_{t}$ to ${\bm{z}}^{src}_{t}$ arises when sampling an audio from ${\bm{z}}^{*}_{T}$ where $\alpha$ is hyper-parameter:
-${\bm{z}}_{t}=\sqrt{\alpha_{t}}{\bm{z}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$ (13)
-A.3 Classifier-free Guidance
-Classifier-free Guidance (CFG) (Ho & Salimans, 2021) is proposed to overcome the limitation of text-conditioned models, where text adherence could be weak.
-However, a higher guidance scale, $\omega$, which is intended to strengthen the model’s fidelity to the text prompt, inadvertently magnifies the accumulated inversion error.
-This becomes problematic in editing scenarios where precise control over the audio synthesis is desired.
-The modified noise estimation in CFG can be expressed as:
-$\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$ (14)
-where $\varnothing$ = (“”) is the embedding of a null text.
-This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{{}^{\prime\prime}}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
-Appendix B Benchmark construction
-B.1 General information
+Acknowledgements.
+The research was supported by the Early Career Scheme (ECS-Hong Kong University of Science and Technology 22201322) and NSFC (No. 62206234) of Mainland China.
+  1. (1)
+    Benchmark construction
+    1. (a)
+      General information
+    2. (b)
+      Annotation Process
+    3. (c)
+      Data Format
+  2. (2)
+    Implementation Details
+    1. (a)
+      Metrics
+  3. (3)
+    Quantitative Results
+  4. (4)
+    Potential Negative Societal Impacts
+  5. (5)
+    Limitations
+  6. (6)
+    Qualitative Results
+    1. (a)
+      Change Instrument
+    2. (b)
+      Add Instrument
+    3. (c)
+      Delete Instrument
+    4. (d)
+      Change Genre
+    5. (e)
+      Change Mood
+    6. (f)
+      Change Rhythm
+    7. (g)
+      Change Background
+    8. (h)
+      Change Melody
+    9. (i)
+      Extract Instrument
+  7. (7)
+    safeguards
+7. Benchmark construction
+7.1. General information
 Here are the details of our ZoME-Bench dataset (Zero-shot Music Editing Benchmark).
@@
 We refactor the original captions to express specific edits and divide them into 10 parts, each representing a different type of editing. A sample and details are shown in the following table 4.
-Editing type id Editing type size origin prompt editing prompt editing instruction
+Editing
+type id
+Editing type size origin prompt editing prompt editing instruction
 0 change instrument 131 ambient acoustic [guitar] music ambient acoustic [violin] music change the instrument from guitar to violin
@@
 9 random 67 / / /
-Table 5: Information of ZoME-Bench dataset
-B.2 Annotation Process
-We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023).
-With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type “change melody" as examples):
+Table 4. Information of ZoME-Bench dataset
+7.2. Annotation Process
+We rebuild our caption from captions for Musiccaps offered by  (agostinelli2023musiclm).
+With the help of ChatGPT-4 (openai2023gpt4), we rebuild the caption with prompt as follows(take type “change melody” as examples):
 Description: “There is a description of a Piece of music, Please judge whether the description has information of melody.
-If not, just answer “Flase", else change its melody properly into the opposite one, just change the adjective and don’t replace any instrument!
-",“blended_word" is [origin melody, changed melody], “emphasize" is [changed melody], “blended_word" and “emphasize" are tuples.
+If not, just answer “Flase”, else change its melody properly into the opposite one, just change the adjective and don’t replace any instrument!
+”,“blended_word” is [origin melody, changed melody], “emphasize” is [changed melody], “blended_word” and “emphasize” are tuples.
 Question: (A mellow, passionate melody from a noisy electric guitar)
-Answer:(“source_prompt": “A mellow, [passionate] melody from a noisy electric guitar", “editing_prompt": “A mellow, [soft] melody from a noisy electric guitar", “blended_word": [“passionate melody", “soft melody"], “emphasize": [“soft melody"])
+Answer:(“source_prompt”: “A mellow, [passionate] melody from a noisy electric guitar”, “editing_prompt”: “A mellow, [soft] melody from a noisy electric guitar”, “blended_word”: [“passionate melody”, “soft melody”], “emphasize”: [“soft melody”])
 Question: (A recording of solo harp music with a dreamy, relaxing melody.)
-Answer: (“source_prompt": “A recording of solo harp music with a dreamy, [relaxing] melody.", “editing_prompt": “A recording of solo harp music with a dreamy, [nervous] melody.", “blended_word": [“relaxing melody",“nervous melody"], “emphasize" :[“nervous melody"])
-Question: (“A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
-Answer: (“source_prompt": “A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", “editing_prompt": “A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", “blended_word": [“harmonized flute melody",“passionate flute melody"]), “emphasize" :[“passionate flute melody"])
-Now we have Question:({origin caption}), Answer(?)"
-In the same way, instructions are appended by prompt as follows (take type “change melody" as examples):
+Answer: (“source_prompt”: “A recording of solo harp music with a dreamy, [relaxing] melody.”, “editing_prompt”: “A recording of solo harp music with a dreamy, [nervous] melody.”, “blended_word”: [“relaxing melody”,“nervous melody”], “emphasize” :[“nervous melody”])
+Question: (“A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions”)
+Answer: (“source_prompt”: “A vintage, emotional song with [passionate] flute melody and soft wooden percussions.”, “editing_prompt”: “A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.”, “blended_word”: [“harmonized flute melody”,“passionate flute melody”]), “emphasize” :[“passionate flute melody”])
+Now we have Question:({origin caption}), Answer(?)”
+In the same way, instructions are appended by prompt as follows (take type “change melody” as examples):
 Description: “There are two descriptions of different pieces of music divided by &, Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
@@
 Answer: change dissonant melody into harmonic melody
-Now we have Question: [‘source prompt’] & [‘editing prompt’], Answer(?)"
+Now we have Question: [‘source prompt’] & [‘editing prompt’], Answer(?)”
 Through this method, supplemented by rounds of manual review, we ensure the quality of this benchmark.
-B.3 Data Format
+7.3. Data Format
 Taking the first piece as an example, we express our data in JSON format with six keys
-"$000$": {
+"000000000000": {
 "editing_prompt": "a live recording of ambient acoustic
@@
 [guitar] music",
-"blended_word": "(\"guitar\", \"violin\")",
-"emphasize": "(\"violin\")",
+"blended_word": "("guitar", "violin")",
+"emphasize": "("violin")",
 "audio_path": "wavs/MusicCaps_-4SYC2YgzL8.wav",
@@
 to violin"
-“Editing_prompt" refers to the edited caption, while “source_prompt" denotes the original caption.
-"Blended_word" indicates the subject to be edited, and “Emphasize" represents the word that should be highlighted.
-“Editing_instruction" provides a description of the editing process.
-Additionally, in the editing type “delete instrument," we introduce another key, “neg_prompt", which helps reduce the likelihood of deleted instruments reappearing.
-Appendix C Implementation Details
-For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023c).
+“Editing_prompt” refers to the edited caption, while “source_prompt” denotes the original caption.
+”Blended_word” indicates the subject to be edited, and “Emphasize” represents the word that should be highlighted.
+“Editing_instruction” provides a description of the editing process.
+Additionally, in the editing type “delete instrument,” we introduce another key, “neg_prompt”, which helps reduce the likelihood of deleted instruments reappearing.
+8. Implementation Details
+For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (liu2023audioldm2).
 Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD.
 These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval.
-In line with the methodology described by  Manor & Michaeli (2024), we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion.
+In line with the methodology described by  manor2024zero, we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion.
 For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12.
-The forward guidance of MEDIC is 1 while the reverse scale is 5.
-We chose these values by exploring different guidance scales, as discussed in Appendix D.
+The forward guidance of MelodyEdit is 1 while the reverse scale is 5.
+We chose these values by exploring different guidance scales, as discussed in Appendix 9.
 We conduct all experiments in NVIDIA 4090.
-For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023c).
-To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD and conduct on an NVIDIA 4090.
-The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
-Our methodology is aligned with the protocol established by  Manor & Michaeli (2024), where we have adopted a forward guidance scale of 3 and a reverse guidance scale of 12 for DDPM inversion.
+Our methodology is aligned with the protocol established by  manor2024zero, where we have adopted a forward guidance scale of 3 and a reverse guidance scale of 12 for DDPM inversion.
 In contrast, the DDIM inversion employs a guidance scale of 5, and SDEdit utilizes a guidance scale of 12.
 For Disentangled Inversion Control, we have determined the forward guidance to be 1 and the reverse scale to be 5.
-These specific guidance scale values are selected after extensive experimental exploration, the details of which are discussed in Appendix D.
-C.1 Metrics
-Objective Metrics There are details about four metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) CLAP Score (Elizalde et al., 2023): This criterion evaluates the degree to which the output conforms to the specified target prompt.
-(2) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio.
-(3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the source audio.
-(4) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
+These specific guidance scale values are selected after extensive experimental exploration, the details of which are discussed in Appendix 9.
+8.1. Metrics
+Objective Metrics There are details about four metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) CLAP Score (elizalde2023clap): This criterion evaluates the degree to which the output conforms to the specified target prompt.
+(2) Struture Distance (ju2023direct): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio.
+(3) LPAPS (iashin2021taming; paissan2023audio): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable), this measure evaluates the consistency of the edited audio with the source audio.
+(4) FAD (Fréchet Audio Distance) (kilgour2018fad): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
+(5) Chroma (Chromagram Similarity) (zhang2024musicmagus): The average cosine similarity between the chromagrams of the original and edited music, which denotes the preservation of pitch contours and rhythm patterns in the music.
+(6) Accuracy (chu2024qwen2): The rate of successful editing judged by the Qwen model, calculated by constructing a question-answer (QA) pair, where the model’s output is compared against the expected result.
+The comparison assesses whether the model has made the correct edit.
 Subjective Metrics To directly reflect the quality of the audio generated, we carry out MOS (Mean Opinion Score) tests.
@@
 Our crowd-sourced subjective evaluation tests were conducted via Amazon Mechanical Turk where participants were paid \$8 hourly.
-Appendix D Quantitative Results
+9. Quantitative Results
 Analyses on Different CFG Scale The lack of systematic experiments that determine the optimal combination of guidance scales for achieving the best editing performance, and analysis of how these guidance scales affect the final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
@@
 1 1 8.56 0.12 1.17 3.25 0.51
-1 2.5 11.97 0.15 2.49 4.54 0.56
+1 2.5 11.97 0.15 2.49 4.54 0.61
 1 5 15.99 0.17 4.22 6.07 0.61
@@
 7.5 7.5 24.16 0.22 7.33 9.17 0.34
-Table 6: Ablation Studies on Different Guidance Scale
+Table 5. Ablation Studies on Different Guidance Scale
 Inference Time We compare the inference time of our method with baselines, and the results are compiled in Table 6.
-MEDIC achieves the comparative inference time with generation models and inversion techniques.
+MelodyEdit achieves the comparative inference time with generation models and inversion techniques.
 We will make an attempt to reduce the inference time of zero-shot music editing in our future work.
@@
 DDPM-Friendly 33.3s
-MEDIC 92.0s
-Table 7: Inference Time across different methods.
-Appendix E Potential Negative Societal Impacts
-MEDIC may also lead to potential negative societal impacts that are worthy of consideration.
-If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MEDIC could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation.
+MelodyEdit 92.0s
+Table 6. Inference Time across different methods.
+10. Potential Negative Societal Impacts
+MelodyEdit may also lead to potential negative societal impacts that are worthy of consideration.
+If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MelodyEdit could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation.
 Hopefully, all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
-Appendix F Limitations
+11. Limitations
 In spite of the remarkable outcome of our method, due to the limitation of the generation model we used, we are incapable of instigating a profound change.
@@
 We will make an attempt to edit more interesting and complex music tasks in the future.
-Appendix G Qualitative Results
-For each type in ZoME-Bench, We provide samples to observe the capability of MEDIC intuitively.
-G.1 Change Instrument
-In Figure 5, we show the capability of MEDIC to change the instrument.
-Here we edit the ground truth music piece with the source prompt “a live recording of ambient acoustic [guitar] music" and editing prompt “a live recording of ambient acoustic [violin] music".
+12. Qualitative Results
+For each type in ZoME-Bench, We provide samples to observe the capability of MelodyEdit intuitively.
+12.1. Change Instrument
+In Figure 7, we show the capability of MelodyEdit to change the instrument.
+Here we edit the ground truth music piece with the source prompt “a live recording of ambient acoustic [guitar] music” and editing prompt “a live recording of ambient acoustic [violin] music”.
 The difference in instruments can be observed in the Mel-spectrum.
@@
 edit: a live recording of ambient acoustic [violin] music
-Figure 5: Editing Type 0 :Change Instrument
-G.2 Add Instrument
-In Figure 6, we show the capability of MEDIC to add more instruments.
-Here we edit the ground truth music piece with the source prompt “a heavy metal instructional audio with a distortion guitar" and editing prompt “a heavy metal instructional audio with a distortion guitar [and drums]".
+Figure 7. Editing Type 0 :Change Instrument
+12.2. Add Instrument
+In Figure 8, we show the capability of MelodyEdit to add more instruments.
+Here we edit the ground truth music piece with the source prompt “a heavy metal instructional audio with a distortion guitar” and editing prompt “a heavy metal instructional audio with a distortion guitar [and drums]”.
 The appearance of the new instrument can be observed in the Mel-spectrum which presents a drum sound of high frequency.
@@
 edit: a heavy metal instructional audio with a distortion guitar [and drums]
-Figure 6: Editing Type 1 Add Instrument
-G.3 Delete Instrument
-In Figure 7, we show the capability of MEDIC to delete instruments.
-Here we edit the ground truth music piece with the source prompt “a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and the editing prompt “a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood".
+Figure 8. Editing Type 1 Add Instrument
+12.3. Delete Instrument
+In Figure 9, we show the capability of MelodyEdit to delete instruments.
+Here we edit the ground truth music piece with the source prompt “a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood” and the editing prompt “a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood”.
 The vanishing of the instrument can be observed in the Mel-spectrum.
@@
 edit: a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood
-Figure 7: Editing Type 2 Delete Instrument
-G.4 Change Genre
-In Figure 8, we show the capability of MEDIC to change the genre of a music piece.
-Here we edit the ground truth music piece with the source prompt “a recording of a solo electric guitar playing [blues] licks" and the editing prompt “a recording of a solo electric guitar playing [rock] licks".
+Figure 9. Editing Type 2 Delete Instrument
+12.4. Change Genre
+In Figure 10, we show the capability of MelodyEdit to change the genre of a music piece.
+Here we edit the ground truth music piece with the source prompt “a recording of a solo electric guitar playing [blues] licks” and the editing prompt “a recording of a solo electric guitar playing [rock] licks”.
 The obvious difference in genre can be observed in the Mel-spectrum.
@@
 edit: a recording of a solo electric guitar playing [rock] licks
-Figure 8: Editing Type 3 Change Genre
-G.5 Change Mood
+Figure 10. Editing Type 3 Change Genre
+12.5. Change Mood
 Mood is an important attribute of music.
-In Figure 9, we show the capability of MEDIC to change the mood of a music piece.
-Here we edit the ground truth music piece with the source prompt “a recording of [aggressive] electronic and video game music with synthesizer arrangements" and editing prompt “a recording of [peaceful] electronic and video game music with synthesizer arrangements".
+In Figure 11, we show the capability of MelodyEdit to change the mood of a music piece.
+Here we edit the ground truth music piece with the source prompt “a recording of [aggressive] electronic and video game music with synthesizer arrangements” and editing prompt “a recording of [peaceful] electronic and video game music with synthesizer arrangements”.
 The change of mood can be observed in the Mel-spectrum.
@@
 edit: a recording of [peaceful] electronic and video game music with synthesizer arrangements
-Figure 9: Editing Type 4 Change Mood
-G.6 Change Rhythm
+Figure 11. Editing Type 4 Change Mood
+12.6. Change Rhythm
 Rhythm represents the speed of the music.
-In Figure 10, we show the capability of MEDIC to change the Rhythm of a music piece.
-Here we edit the ground truth music piece with the source prompt “a [slow] tempo ukelele tuning recording with static" and the editing prompt “a [fast] tempo ukelele tuning recording with static".
+In Figure 12, we show the capability of MelodyEdit to change the Rhythm of a music piece.
+Here we edit the ground truth music piece with the source prompt “a [slow] tempo ukelele tuning recording with static” and the editing prompt “a [fast] tempo ukelele tuning recording with static”.
 The change of Rhythm can be observed in the Mel-spectrum.
@@
 edit: a [fast] tempo ukelele tuning recording with static
-Figure 10: Editing Type 5 Change Rhythm
-G.7 Change Background
-In Figure 11, we show the capability of MEDIC to change the background of the instrument of a music piece.
-Here we edit the ground truth music piece with the source prompt “an amateur ukulele recording with a [medium to uptempo] pace" and editing prompt “an amateur ukulele recording with a [steady and rhythmic] pace".
+Figure 12. Editing Type 5 Change Rhythm
+12.7. Change Background
+In Figure 13, we show the capability of MelodyEdit to change the background of the instrument of a music piece.
+Here we edit the ground truth music piece with the source prompt “an amateur ukulele recording with a [medium to uptempo] pace” and editing prompt “an amateur ukulele recording with a [steady and rhythmic] pace”.
 The change of instrument background can be observed in the Mel-spectrum.
-source: an amateur ukulele recording with a [medium to uptempo] pace" and editing prompt
+source: an amateur ukulele recording with a [medium to uptempo] pace” and editing prompt
 edit: an amateur ukulele recording with a [steady and rhythmic] pace
-Figure 11: Editing Type 6 Change Background
-G.8 Change Melody
-In Figure 12, we show the capability of MEDIC to change the melody of a music piece.
-Here we edit the ground truth music piece with the source prompt “this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo" and editing prompt “this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo".
+Figure 13. Editing Type 6 Change Background
+12.8. Change Melody
+In Figure 14, we show the capability of MelodyEdit to change the melody of a music piece.
+Here we edit the ground truth music piece with the source prompt “this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo” and editing prompt “this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo”.
 The change of Melody can be observed in the Mel-spectrum.
@@
 edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
-Figure 12: Editing Type 7 Change Melody
-G.9 Extract Instrument
-In Figure 13, we show the capability of MEDIC to extract one certain instrument of a music piece.
-Here we edit the ground truth music piece with the source prompt “a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]" and editing prompt “a reggae rhythm recording with bongos".
+Figure 14. Editing Type 7 Change Melody
+12.9. Extract Instrument
+In Figure 15, we show the capability of MelodyEdit to extract one certain instrument of a music piece.
+Here we edit the ground truth music piece with the source prompt “a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]” and editing prompt “a reggae rhythm recording with bongos”.
 The change of instruments can be observed in the Mel-spectrum.
@@
 edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
-Figure 13: Editing Type 8 Extract Instrument
-Appendix H safeguards
+Figure 15. Editing Type 8 Extract Instrument
+13. safeguards
 In the processing of the data and models involved in this study, we fully considered the potential risks.